### v2

# EXTENSIONS OF UNIRATIONAL GROUPS

Abstract
We undertake a study of extensions of unirational algebraic groups. We prove that extensions of unirational groups are also unirational over fields of degree of imperfection $1$, but that this fails over every field of higher degree of imperfection, answering a question of Achet. We also initiate a study of those groups which admit filtrations with unirational graded pieces, and show that one may deduce unirationality of unipotent groups from unirationality of certain quotients.
1 Introduction
In [Ach1, Question 4.9], Achet posed the following question: Is a commutative extension of two unirational algebraic $K$-groups necessarily also unirational? His motivation lies in the study of the maximal unirational $K$-subgroup $G_{\mathrm{uni}}$ of an algebraic group, as an affirmative answer implies that, for commutative $G$, the group $G/G_{\mathrm{uni}}$ admits no nontrivial unirational subgroup. The question trivially admits an affirmative answer when $K$ is perfect, as in this case all connected linear algebraic $K$-groups are unirational [Bor, Ch. V,Th. 18.2(ii)].
In this paper, we answer Achet’s question, and more generally undertake a study of extensions of unirational algebraic groups. In §2, we prove that every extension of unirational groups is unirational over fields of degree of imperfection $1$ (Theorem 2.4). However, this fails over every field of degree of imperfection >1>1. In fact, in §3 we construct over every such field a commutative $p$-torsion extension of unirational wound unipotent groups which is not unirational (Example 3.4). We also construct a non-unirational commutative $p$-torsion wound unipotent extension of the additive group by a unirational wound unipotent group (Example 3.5). Thus Achet’s question has a negative answer over every field of degree of imperfection >1>1.
This naturally leads one to introduce a new class of algebraic groups: those which admit a filtration whose successive quotients are unirational. We call such groups ext-unirational, and in §4 we show that ext-unirationality of a connected linear algebraic $G$ over a field of characteristic $p$ is equivalent to ext-unirationality of the maximal commutative $p$-torsion quotient $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}$ (Proposition 4.3), as well as to that of the centralizer $Z_{G}(T)$ of any torus $T$ of $G$ (Proposition 4.6).
Finally, in §5, we continue with the theme of deducing unirationality properties from quotients, but for unirationality itself rather than ext-unirationality. More precisely, we prove that over a field $K$ of degree of imperfection $r$, a smooth connected unipotent $K$-group $U$ is unirational if and only if the quotient $U/\mathscr{D}_{r}U$ of $U$ by the $r$th term in its lower central series is (Theorem 5.7). Of particular note is the case $r\leq 2$, which tells us that, when $K$ has degree of imperfection $\leq 2$, unirationality of $U$ is equivalent to that of its maximal commutative quotient $U^{\operatorname{\rm{ab}}}$. We show that this is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent $K$-group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not (Example 5.9).
Notation and Conventions
Throughout this paper, $K$ denotes a field, and when it appears, $p$ denotes a prime number equal to the characteristic of $K$. By a linear algebraic $K$-group, or a linear algebraic group over $K$, we shall mean a smooth affine $K$-group scheme. Recall that the degree of imperfection of a field $K$ of characteristic $p$ is defined to be $\mathrm{log}_{p}[K:K^{p}]$. This quantity is always either infinite or a nonnegative integer. The degree of imperfection of $K$ also equals $\mathrm{dim}_{K}(\Omega^{1}_{K/\mathbf{F}_{p}})$ as well as the maximal size of a $p$-independent subset of $K$ [Mat, Th. 26.5].
2 Degree of imperfection 11
In this section we will prove that an arbitrary (not necessarily commutative) extension of unirational algebraic $K$-groups must be unirational when $K$ has degree of imperfection $1$. We begin by tying up a loose end concerning a basic permanence property of permawound unipotent groups. Recall that a smooth unipotent $K$-group scheme $U$ is called permawound when, for any exact sequence of finite type $K$-group schemes
$U\longrightarrow E\longrightarrow{\mathbf{G}}_{\rm{a}}\longrightarrow 1,$
$E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$.
Proposition 2.1.
Given an exact sequence
$1\longrightarrow U^{\prime}\longrightarrow U\longrightarrow U^{\prime\prime}\longrightarrow 1$
of unipotent $K$-groups with $U^{\prime},U^{\prime\prime}$ permawound, $U$ is also permawound.
Proof.
If $K$ is perfect, then the assertion is trivially true by [Ros2, Prop. 5.2(i)], so assume that $K$ is imperfect. Then $U^{\prime}$ and $U^{\prime\prime}$ are connected [Ros2, Prop. 6.2], hence so is $U$. We proceed by dimension induction. Note first that, since every quotient of $U$ is an extension of a quotient of $U^{\prime\prime}$ by a quotient of $U^{\prime}$, and since permawoundness is (by definition) inherited by quotients, it follows that we may assume that every quotient of $U$ by a positive-dimensional normal $K$-subgroup is permawound. Consider first the case in which $U$ is not wound. Suppose that we have an exact sequence of finite type $K$-groups
[…]
We must show that $E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$. If $f$ does not kill the maximal split $K$-subgroup $U_{s}$ of $U$, then this is immediate. Otherwise, $f(U)$ is a quotient of $U/U_{s}$, hence permawound, so $E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$ in this case as well. Hence $U$ is permawound.
Now suppose that $U$ is wound. If $U^{\prime}$ is trivial, the assertion is immediate, so assume that $U^{\prime}\neq 1$. Then we claim that there is a nontrivial smooth connected $K$-subgroup $V\subset U^{\prime}$ that is central in $U$. Indeed, define a sequence of subgroups $V_{n}\subset U^{\prime}$ by the following recursive formula: $V_{0}:=U^{\prime}$, and $V_{n+1}:=[U,V_{n}]$ for $n\geq 0$. Then one verifies by induction that the $V_{n}$ are normal subgroups of $U$, and that they form a descending sequence. Because $U$ is unipotent, it is nilpotent, hence $V_{n+1}=0$ for some $n\geq 0$. If we choose $n$ to be minimal, then $V:=V_{n}\subset U^{\prime}$ is nontrivial and central in $U$, which proves the claim.
Let $\overline{U}:=U/V$, let $\overline{U}_{s}\subset\overline{U}$ denote the maximal split subgroup, and let $\overline{U}_{w}:=\overline{U}/\overline{U}_{s}$ denote the wound quotient, which is permawound because $V$ is nontrivial. Because $V\subset U$ is central, the commutator map $U\times U\rightarrow U$ descends to a map $c\colon\overline{U}\times\overline{U}\rightarrow U$. We claim that $c$ further descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$. Indeed, let $u_{1},u_{2}\in\overline{U}(K_{s})$. Then $c$ restricts to a map $u_{1}\overline{U}_{s}\times u_{2}\overline{U}_{s}\rightarrow U_{K_{s}}$ which must be constant because $\overline{U}_{s}^{2}$ is split while $U_{K_{s}}$ is wound. Because $\overline{U}^{2}(K_{s})$ is Zariski dense in $\overline{U}^{2}$, it follows that $c$ is $\overline{U}_{s}$-invariant in both arguments, hence $c$ indeed descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$ which we by abuse of notation also denote by $c$. Because $\overline{U}_{w}$ is wound and permawound, it is commutative [Ros3, Cor. 10.4], so $c$ is the constant map to $1\in U(K)$ by [Ros3, Lem. 10.2]. It follows that $U$ is commutative. Because commutative permawound groups are trivially weakly permawound [Ros2, Def. 5.1], $U^{\prime}$ and $U^{\prime\prime}$ are weakly permawound, hence so is $U$ [Ros2, Prop. 5.6]. It follows that $U$ is permawound [Ros2, Prop. 6.9]. ∎
Next we verify that, for fields of degree of imperfection $1$, unirationality for unipotent groups is the same as permawoundness.
Proposition 2.2.
Let $K$ be a field of degree of imperfection $1$. A unipotent $K$-group $U$ is unirational if and only if it is permawound.
Proof.
For the if direction, use [Ros3, Th. 1.9(i)]. For the converse, [Ros3, Prop. 9.6] yields the result when $U$ is wound. For arbitrary $U$, unirationality implies that $U$ is smooth and connected. Then we have an exact sequence
$1\longrightarrow U_{s}\longrightarrow U\longrightarrow U_{w}\longrightarrow 1$
with $U_{s}$ split and $U_{w}$ wound. Because $U$ is unirational, so is $U_{w}$, which is therefore permawound. Then [Ros2, Prop. 5.3] and Proposition 2.1 show that $U$ is permawound. ∎
We require the following simple lemma.
Lemma 2.3.
Given an extension of finite type $K$-group schemes
$1\longrightarrow T\longrightarrow G\longrightarrow U\longrightarrow 1$
with $T$ a torus and $U$ a unirational $K$-group scheme, $G$ is also unirational.
Proof.
We may assume that $K$ is separably closed [Ros3, Th. 1.6]. Let $f\colon X\rightarrow U$ be a dominant map from a dense open subscheme $X\subset\mathbf{A}^{n}$. Then $G$ is in particular a $T$-torsor over $U$, and $f^{*}(G)$ is a $T$-torsor over $X$, which is trivial because $T$ is split and $\operatorname{Pic}(X)=0$. Thus $f^{*}(G)$ is rational, so the dominant map $f^{*}(G)\rightarrow G$ proves that $G$ is unirational. ∎
Theorem 2.4.
Let $K$ be a field of degree of imperfection $1$, and suppose given an extension
[…]
of finite type $K$-group schemes. If $G^{\prime}$ and $G^{\prime\prime}$ are unirational, then so is $G$.
Proof.
First, unirational $K$-schemes are necessarily generically smooth and geometrically connected, so $^{\prime}, G′′G^{\prime\prime}$ are smooth and connected. They are also linear algebraic, for if not, then – at least over $\overline{K}$ – they would admit nonzero abelian variety quotients, which would then be unirational. But a nonzero abelian variety is never unirational. (Any map from a nonempty open subscheme of $\mathbf{P}^{1}$ into an abelian variety $A$ extends over $\mathbf{P}^{1}$, and abelian varieties do not admit nonconstant maps from $\mathbf{P}^{1}$.) It follows that $G$ is also a connected linear algebraic group.
Choose a maximal $K$-torus $T\subset G$, and let $Z_{G}(T)$ denote the centralizer in $G$ of $T$ (and similarly with the other centralizers below). Let $T^{\prime\prime}:=\pi(T)$, a maximal $K$-torus of $G^{\prime\prime}$ [Bor, Ch. IV, Prop. 11.14], and let $T^{\prime}:=T\cap G^{\prime}$, a maximal $K$-torus of $G^{\prime}$ [CGP, Cor. A.2.7]. Then we have an exact sequence
$1\longrightarrow Z_{G^{\prime}}(T^{\prime})/T^{\prime}\longrightarrow Z_{G}(T)/T\longrightarrow Z_{G^{\prime\prime}}(T^{\prime\prime})/T^{\prime\prime}\longrightarrow 1,$ (2.1)
with exactness on the right using [Bor, Ch. IV,11.14,Cor. 2]. The groups $Z_{G^{\prime\prime}}(T^{\prime\prime})$ and $Z_{G^{\prime}}(T^{\prime})$ are unirational [Ros3, Prop 7.12], hence so are the two groups on the end in (2.1). They are also unipotent, the unipotence a consequence of the maximality of the respective tori combined with [Bor, Ch. IV,11.5,Cor. 2]. Now an extension of unirational unipotent groups over a field of degree of imperfection $1$ is still unirational by combining Propositions 2.2 and 2.1. Thus $Z_{G}(T)/T$ is unirational, hence so is $Z_{G}(T)$ (Lemma 2.3), hence so is $G$ [Ros3, Prop. 7.12]. ∎
3 Examples of extensions of unirational groups that are not unirational
In this section we will show by example – over every field of degree of imperfection >1>1 – that unirationality is not inherited by extensions in general. Our examples will be of two main types: Our first example will be an extension of a wound unirational group by a wound unirational group, while our second will be an extension of ${\mathbf{G}}_{\rm{a}}$ by a wound unirational group. Of course, any extension of a unirational group by a split unipotent one is unirational, as ${\rm{H}}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$. Our examples will all be commutative and $p$-torsion, hence they provide a negative answer to Achet’s question over fields of degree of imperfection >1>1. The proofs that these examples work will require various auxiliary lemmas, to which we now turn.
For $\alpha\in K$ and $n\geq 1$, we define the $K$-group
$V_{n,\alpha}:=\left\{-S+\alpha^{p-1}S^{p}+\sum_{\begin{subarray}{c}0\leq j<p^{n}\\ j\not\equiv-1\pmod{p}\end{subarray}}\alpha^{j}S_{j}^{p^{n}}=0\right\}\subset{\mathbf{G}}_{\rm{a}}^{p^{n}-p^{n-1}+1}.$ (3.1)
The next lemma gives a different description of this group. Before stating it, let us recall some basic notions regarding $p$-polynomials. An element of $K[X_{1},\dots,X_{n}]$ is called a $p$-polynomial when it is a sum of terms of the form $cX_{i}^{p^{r}}$ with $c\in K$ and $r\geq 0$. The sum $P$ of the leading terms in each variable $X_{i}$ is called the principal part of $F$, and we say that $F$ is reduced when $P$ has no nontrivial zeroes over $K$: If $P(x_{1},\dots,x_{n})=0$ with $x_{i}\in K$, then $x_{i}=0$ for all $i$. We say that $F$ is universal when the homomorphism that it induces $K^{n}\rightarrow K$ is surjective.
Lemma 3.1.
If $\alpha\in K-K^{p}$, then
$V_{n,\alpha}\simeq\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}\left[\operatorname{R}_{K(\alpha^{1/p^{n}})/{K(\alpha^{1/p^{n-1}})}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}\right],$
where we regard ${\mathbf{G}}_{\rm{m}}$ as sitting inside $\operatorname{R}_{K(\alpha^{1/p^{n}})/{K(\alpha^{1/p^{n-1}})}}({\mathbf{G}}_{\rm{m}})$ in the natural manner, via the inclusion $A^{\times}\subset(A\otimes_{K(\alpha^{1/p^{n-1}})}K(\alpha^{1/p^{n}}))^{\times}$ for every $K(\alpha^{1/p^{n-1}})$-algebra $A$.
Note in particular that $V_{n,\alpha}$ is unirational. This also may be seen as folllows: $V_{n,\alpha}$ is defined over the field $\mathbf{F}_{p}(\alpha)$, so we may verify its unirationality over this field, which has degree of imperfection $1$. But over $\mathbf{F}_{p}(\alpha)$, $V_{n,\alpha}$ is permawound [Ros2, Props. 6.4,6.9,3.5], hence unirational [Ros3, Th. 1.9(i)].
Proof.
We first note that all constructions above are valid over $\mathbf{F}_{p}(\alpha)\subset K$, and commute with arbitrary separable extension on $K$. Thus we may assume that $K=\mathbf{F}_{p}(\alpha)$, and in particular that $K$ has degree of imperfection $1$. For ease of notation, let us denote $V_{n,\alpha}$ by $V$, and $K(\alpha^{1/p^{n}})=K^{1/p^{n}}$ by $K_{n}$. We claim that one has an isomorphism of $K_{n-1}$-groups
$V_{K_{n-1}}\simeq{\mathbf{G}}_{\rm{a}}^{(p-1)(p^{n-1}-1)}\times W,$ (3.2)
where
$W\simeq\left\{-Y+\alpha^{p-1}Y^{p}+\sum_{0\leq j<p-1}\alpha^{j}Y_{j}^{p^{n}}=0\right\}\subset{\mathbf{G}}_{\rm{a}}^{p}.$ (3.3)
Indeed, over $K_{n-1}$ one may apply the invertible change of variables which fixes $S$ and $S_{j}$ for i>p1i>p-1, and such that
$S_{j}\mapsto S_{j}-\sum_{\begin{subarray}{c}j<i<p^{n}\\ i\equiv j\pmod{p}\end{subarray}}\alpha^{(i-j)/p^{n}}S_{i},\hskip 21.68121pt0\leq j<p-1.$
This has the effect of replacing the equation for $V$ to that of $W$ (after renaming the $S$ variables as $Y$ variables), and one has a copy of ${\mathbf{G}}_{\rm{a}}$ for each of the additional variables $S_{j}$, j>p1j>p-1, not appearing in the equation, which proves (3.2).
We further massage the equation for $W$ by making the invertible change of variables over $K_{n-1}$ which fixes $Y_{j}$ and does the following to $Y$:
$Y\mapsto\alpha^{\frac{1-p^{n-1}}{p^{n-1}}}Y-\sum_{\begin{subarray}{c}0\leq j<p-1\\ 1\leq i<n\end{subarray}}\alpha^{\frac{j+1-p^{n-i}}{p^{n-i}}}Y_{j}^{p^{i}}.$
One verifies that this has the effect (after multiplying through by $\alpha^{\frac{p^{n-1}-1}{p^{n-1}}}$) of replacing the equation (3.3) for $W$ with the following equation:
$W\simeq\left\{-Y+\alpha^{\frac{p-1}{p^{n-1}}}Y^{p}+\sum_{0\leq j<p-1}\alpha^{\frac{j}{p^{n-1}}}Y_{j}^{p}=0\right\}.$ (3.4)
By [Oes, Ch. VI,Prop. 5.3], this last equation describes the $K_{n-1}$-group $\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, so via (3.2), we obtain a nonzero $K_{n-1}$-homomorphism $V_{K_{n-1}}\rightarrow\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, whence a nonzero homomorphism $g\colon V\rightarrow\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ over $K$. It remains to prove that this (and indeed, any such) homomorphism is an isomorphism. The principal part of the polynomial (3.1) defining $V$ is reduced and universal over $F$ (for instance, using [Ros2, Prop. 3.5(ii)]), so $V$ is permawound over $F$ [Ros2, Props. 6.4,6.9], hence unirational [Ros3, Th. 1.9(i)]. Thus $\operatorname{im}(g)\subset\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is a nonzero unirational $K$-subgroup. But the latter group is a minimal unirational $K$-group in the sense that its only nonzero unirational $K$-subgroup is itself [Ach2, Props. 2.18,2.20], hence $g$ is surjective. To prove that $g$ is injective, we note that dimension considerations imply that $\ker(g)$ is finite. Because $V/\ker(g)\simeq\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is wound unipotent, $\ker(g)$ is weakly permawound [Ros2, Prop. 5.5], hence has strictly positive dimension if nontrivial (for instance, by [Ros2, Th. 9.5]; recall that weak permawoundness is insensitive to algebraic separable extension of the ground field [Ros2, Prop. 6.7]). Thus $\ker(g)=0$, so $g$ is an isomorphism. ∎
Lemma 3.2.
Suppose given a separably closed field $K$ and a commutative $p$-torsion unirational wound unipotent $K$-group scheme $U$. Then $U$ is generated by maps of the form $V_{n,\alpha}\rightarrow U$ with $n\geq 1$ and $\alpha\in K-K^{p}$.
Proof.
Because $U$ is unirational, it is generated by maps from nonempty open subschemes of $\mathbf{P}^{1}$ with a rational point lying above the identity. But any such map $X\rightarrow U$ factors through a homomorphism $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ for some finite closed subscheme $D\subset\mathbf{P}^{1}$ by [BLR, §10.3, Thm. 2]. (For a more detailed explanation of how this follows from that theorem, see the discussion before Theorem 2.2 in [Ros1, §2].) Because $U$ is wound, $D$ may be taken to be reduced by [Ros1, Lem. 2.3]. Write $D=\coprod\{x_{i}\}$ with the $x_{i}$ (reduced) closed points in $\mathbf{P}^{1}$. Then $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{i}\operatorname{R}_{K(x_{i})/K}({\mathbf{G}}_{\rm{m}})$. Hence $U$ is generated by maps from $\operatorname{R}_{K(x)/K}({\mathbf{G}}_{\rm{m}})$ as $x$ varies over closed points of $\mathbf{P}^{1}$. Because $K$ is separably closed, $K(x)/K$ is a (finite) purely inseparable extension, and it is a primitive extension (generated by a single element) by [Ros3, Lem. 5.4]. We therefore conclude that $U$ is generated by maps from $\operatorname{R}_{L/K}({\mathbf{G}}_{\rm{m}})$ as $L$ varies over primitive nontrivial purely inseparable extensions of $K$. But any such extension is of the form $K(\alpha^{1/p^{n}})$ for some $\alpha\in K-K^{p}$ and $n\geq 1$. The multiplication by $p$ map on $\operatorname{R}_{K(\alpha^{1/p^{n}})/K}({\mathbf{G}}_{\rm{m}})$ has image $\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}({\mathbf{G}}_{\rm{m}})$, as may, for instance, be checked on $\mathbf{F}_{p}(\alpha)_{s}$-points, since both groups are smooth and defined over $\mathbf{F}_{p}(\alpha)$. Thus, because $U$ is $p$-torsion, it is generated by maps from groups of the form
$\operatorname{R}_{K(\alpha^{1/p^{n}})/K}({\mathbf{G}}_{\rm{m}})/\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}({\mathbf{G}}_{\rm{m}})\simeq\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}\left[\operatorname{R}_{K(\alpha^{1/p^{n}})/{K(\alpha^{1/p^{n-1}})}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}\right],$
the last equality by [CGP, Cor. A.5.4(3)]. Now apply Lemma 3.1. ∎
Lemma 3.3.
Let $\alpha,\lambda\in K-K^{p}$, and let $f\colon V_{n,\alpha}\rightarrow V_{1,\lambda}\subset{\mathbf{G}}_{\rm{a}}^{p}$ be a nonzero $K$-group homomorphism. Then $K^{p}(\alpha)=K^{p}(\lambda)$, and $f$ is ((in each coordinate)) homogeneous linear in $S$, and homogeneous of degree $p^{n-1}$ in the $S_{j}$. Further, using the variables $X,X_{i}$ on $V_{1,\lambda}$ ((instead of $S,S_{j}$)), the $X$ coordinate is of the form $aS$ with $0\neq a\in K$.
Proof.
Because $V_{1,\lambda}$ is a minimal unirational $K$-group – in the sense that its only nonzero unirational $K$-subgroup is itself [Ach2, Props. 2.18,2.20]$f$ is surjective. The group $V_{n,\alpha}$ splits over $K(\alpha^{1/p^{\infty}})$, hence so does $V_{1,\lambda}$, so we deduce that $\lambda^{1/p}\in K(\alpha^{1/p^{\infty}})$. It follows that $K^{p}(\alpha)=K^{p}(\lambda)$. Let $\{\alpha\}\coprod\mathscr{B}$ be a $p$-basis for $K$. Then the extension $L:=K(\mathscr{B}^{p^{-\infty}})$ obtained by adjoining all $p$-power roots of elements of $\mathscr{B}$ to $K$ has $\{\alpha\}$ as a $p$-basis. Further, one has $\alpha\in L^{p}(\lambda)$, so $\lambda\notin L^{p}$ and all hypotheses are preserved upon scalar extension to $L$. We may therefore assume that $K$ has degree of imperfection $1$.
We claim that we have for $m\geq 1$ a surjection $f_{m}\colon V_{m+1,\alpha}\rightarrow V_{m,\alpha}$ with kernel a power of $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$. Indeed, using the equation (3.1), if we use the letter $S$ to denote the variables for $V_{m+1,\alpha}$, and the letter $T$ to denote those for $V_{m,\alpha}$, we have the map $(S,(S_{j})_{j})\mapsto(Z,(Z_{j})_{j})$, where $Z:=S$ and $Z_{j}:=\sum_{i=0}^{p-1}\alpha^{i}S_{j+p^{m}i}^{p}$. The kernel of this map is a power of the $K$-group
$W:=\left\{\sum_{i=0}^{p-1}\alpha^{i}X_{i}^{p}=0\right\}\subset{\mathbf{G}}_{\rm{a}}^{p},$
which is isomorphic to $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ [Ros2, Prop. 7.4].
By [Ros2, Lem. 7.9], therefore, and an easy induction, any map $V_{n,\alpha}\rightarrow V_{1,\lambda}$ must factor through the map $V_{n,\alpha}\rightarrow V_{1,\alpha}$ obtained by iterating the maps $f_{m}$ above. The lemma therefore reduces to the case $n=1$, where we must show that the map in question is obtained by a linear change of variables, and that the coefficient of $S$ in $X$ is nonzero. (That $X$ may be written uniquely as a $p$-polynomial of degree $\leq 1$ in $S$ follows from [Ros1, Prop. 6.4].) For this, we may extend scalars to $K_{s}$ and thereby assume that $K$ is separably closed. The lemma then follows from [Ros2, Lem. 7.1,Prop. 9.7], since, for a linear change of variables that transforms the polynomial $X+\lambda^{p-1}X+\sum_{i=0}^{p-2}\lambda^{i}S_{i}^{p}$ into a nonzero scalar multiple of $S+\alpha^{p-1}S+\sum_{i=0}^{p-2}\alpha^{i}S_{i}^{p}$, $X=X(S,S_{i})$ must be a nonzero multiple of $S$. ∎
Now we give our examples.
Example 3.4.
Let $K$ be a field of degree of imperfection >1>1, and let $\lambda,\mu\in K$ be $p$-independent. (This has various equivalent definitions. For our purposes, the reader can take it to mean that $K^{p}(\lambda,\mu):K^{p}]=p^{2}$.) We use the equation (3.1) for the variables on $V_{1,\lambda}$, using the letter $Y$ instead of $S$ to denote the variables: $Y,Y_{0},\dots,Y_{p-2}$. We have an exact sequence
[…]
where
$F_{\mu}(X_{0},\dots,X_{p-1}):=-X_{p-1}+\sum_{i=0}^{p-1}\mu^{i}X_{i}^{p}.$
Consider the homomorphism $f\colon V_{1,\lambda}\rightarrow{\mathbf{G}}_{\rm{a}}$ defined by $Y$, and let $E\in\operatorname{Ext}^{1}(V_{1,\lambda},V_{1,\mu})$ be its image under the connecting map $\delta$ associated to (3.5). Then we claim that $E$ is not unirational, despite the fact that $V_{1,\lambda}$ and $V_{1,\mu}$ are.
In proving that $E$ is not unirational, we may assume that $K$ is separably closed. Suppose that $E$ is unirational, and we will derive a contradiction. The group $E$ is commutative and $p$-torsion, hence by Lemma 3.2, there exist $\alpha\in K-K^{p}$, $n\geq 1$, and a homomorphism $V_{n,\alpha}\rightarrow E$ whose image in $V_{1,\lambda}$ is nonzero. Stated differently, there is a nonzero homomorphism $g\colon V_{n,\alpha}\rightarrow V_{1,\lambda}$ such that $\delta(f\circ g)=0$, or equivalently, $f\circ g\colon V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}$ is of the form $F_{\mu}(h)$ for some homomorphism $h\colon V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}^{p}$.
By Lemma 3.3 (and in the notation of that lemma), $K^{p}(\alpha)=K^{p}(\lambda)$, and
$f\circ g=cS\mbox{ for some }0\neq c\in K.$ (3.6)
Applying [Ros1, Prop. 6.4], there exist $p$-polynomials $X_{0},\dots,X_{p-1}\in K[S,(S_{j})_{j}]$ of degree $\leq 1$ in $S$ such that
$f\circ g=cS=-X_{p-1}+\sum_{i=0}^{p-1}\mu^{i}X_{i}^{p}$ (3.7)
as homomorphisms $V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}$. Applying [Ros1, Prop. 6.4] once more, if we use the equation (3.1) to eliminate $S^{p}$ from the $X_{i}^{p}$ terms above, then this becomes an identity of polynomials in $S$ and the $S_{j}$. Write
$X_{i}=r_{i}S+\sum_{j}r_{ij}S_{j}^{p^{n-1}}+G_{i}$
with $r_{i},r_{ij}\in K$, and where $G_{i}$ does not involve either $S$ or $S_{j}^{p^{n-1}}$. We claim that $G_{i}$ has degree $p^{n-1}$. Indeed, if not, then looking at the leading term in some $S_{j}$ in (3.7) would yield a nonzero solution over $K$ to the equation $\sum_{i=0}^{p-1}\mu^{i}x_{i}^{p}=0$, which would violate the fact that $\mu\notin K^{p}$.
Comparing coefficients of $S_{0}^{p^{n}}$ in (3.7) now yields
$\alpha^{1-p}\sum_{i=0}^{p-1}\mu^{i}r_{i}^{p}=\sum_{i=0}^{p-1}\mu^{i}r_{i0}^{p}.$
If some $r_{i}$ is nonzero, then $\alpha\in K^{p}(\mu)$. Since $\alpha\in K^{p}(\lambda)$, the $p$-independence of $\lambda,\mu$ would then imply that $\alpha\in K^{p}$, which is false. Thus $r_{i}=0$ for all $0\leq i<p$. It follows that each $X_{i}$ is a $p$-polynomial in the $S_{j}$ only, hence by (3.7), $c=0$, in violation of (3.6). This contradiction shows that $E$ is not unirational.
Example 3.5.
Let $K$ be a field of degree of imperfection >1>1, and let $\lambda,\mu\in K$ be $p$-independent elements. Assume first that $p>2$. We will sketch how to modify the example and argument below when $p=2$. Define $U\subset{\mathbf{G}}_{\rm{a}}^{p+1}$ to be the $K$-group scheme
$U:=\left\{-X_{p-1}+\mu Y^{p}+\sum_{i=0}^{p-1}\lambda^{i}X_{i}^{p}=0\right\}.$ (3.8)
The map $(Y,X_{0},\dots X_{p-1})\mapsto Y$ defines a surjective homomorphism $U\rightarrow{\mathbf{G}}_{\rm{a}}$ with kernel $V_{1,\lambda}$. Thus $U$ is an extension of ${\mathbf{G}}_{\rm{a}}$ by the unirational group $V_{1,\lambda}$, but we claim that $U$ is not unirational.
We first note that $K/\mathbf{F}_{p}(\lambda,\mu)$ is separable, precisely because $\lambda,\mu$ are $p$-independent over $K$ [Mat, Th. 26.6]. Thus the non-unirationality of $U$ may be checked over $\mathbf{F}_{p}(\lambda,\mu)$ [Ros3, Th. 1.6], hence we may assume that $\lambda,\mu$ form a $p$-basis for $K$. We may also extend scalars to $K_{s}$ and thereby additionally assume that $K$ is separably closed. Assume for the sake of contradiction that $U$ is unirational. Then by Lemma 3.2, there exist $n>0$,
$\alpha\in K-K^{p},$ (3.9)
and a homomorphism $g\colon V_{n,\alpha}\rightarrow U$ whose projection onto the $Y$ coordinate is nonzero. By [Ros1, Prop. 6.4], $g=(Y(S,(S_{j})_{j}),X_{0}(S,(S_{j})_{j}),\dots,X_{p-1}(S,(S_{j})_{j}))$, where
$Y$ $=c\cdot S+F((S_{j})_{j})$
$X_{i}$ $=c_{i}\cdot S+F_{i}((S_{j})_{j})$ (3.10)
for some $c,c_{i}\in K$ and some $p$-polynomials $F,F_{i}$, and
$Y(S,(S_{j})_{j})\neq 0.$ (3.11)
These expressions must satisfy the defining equation (3.8) for $U$.
Substituting the expressions (3.5) into (3.8), and using the equation (3.1) for $V_{n,\alpha}$ to eliminate $S^{p}$, we obtain
$c_{p-1}S+\alpha^{1-p}\left(\mu c^{p}+\sum_{i=0}^{p-1}\lambda^{i}c_{i}^{p}\right)\left[S-\sum_{\begin{subarray}{c}0\leq j<p^{n}\\ j\not\equiv-1\pmod{p}\end{subarray}}\alpha^{j}S_{j}^{p^{n}}\right]-F_{p-1}+\mu F^{p}+\sum_{i=0}^{p-1}\lambda^{i}F_{i}^{p}=0.$ (3.12)
Because the above expression has degree <p<p in $S$, it is an identity of $p$-polynomials in the variables $S,S_{j}$ [Ros1, Prop. 6.4]. We claim that
$\mathrm{deg}(F_{p-1})\leq p^{n-1}.$ (3.13)
Indeed, if not, then we have $p^{d}:=\mathrm{deg}(F_{p-1})>p^{n-1}$. Let $\beta_{ij}$ be the coefficient of $S_{j}^{p^{d}}$ in $F_{i}$, and let $\beta_{j}$ be the coefficient of $S_{j}^{p^{d}}$ in $F$, so $\beta_{p-1,j}\neq 0$ for some $j$. Then comparing coefficients of $S_{j}^{p^{d+1}}$ in (3.12) yields $\mu\beta_{j}^{p}+\sum_{i=0}^{p-1}\lambda^{i}\beta_{ij}^{p}=0$, and the $p$-independence of $\lambda,\mu$ then shows that all $\beta_{ij}=0$, in violation of the nonvanishing of $\beta_{p-1,j}$. Thus (3.13) holds.
Let $r_{0}:=\mu c^{p}+\sum_{i=0}^{p-1}\lambda^{i}c_{i}^{p}$. We claim that $r_{0}\neq 0$. Indeed, suppose that $r_{0}=0$ and we will obtain a contradiction. We have $c=0$ by $p$-independence of $\lambda,\mu$, hence $F\neq 0$ by (3.11) and (3.5). Additionally, (3.12) would then yield
$c_{p-1}S-F_{p-1}+\mu F^{p}+\sum_{i=0}^{p-1}\lambda^{i}F_{i}^{p}\stackrel{{\scriptstyle?}}{{=}}0.$ (3.14)
Let $p^{d}$ denote the maximum degree in of $F,F_{i}$ (well-defined since $F\neq 0$), and suppose that one of $F,F_{i}$ involves a term $S_{j}^{p^{d}}$ with nonzero coefficient. Then comparing the coefficients of $S_{j}^{p^{d+1}}$ in (3.14) yields a nontrivial solution over $K$ to the equation $\mu a^{p}+\sum_{i=0}^{p-1}\lambda^{i}a_{i}^{p}=0$, in violation of the $p$-independence of $\lambda,\mu$. This contradiction shows that indeed $r_{0}\neq 0$.
Let
$R:=\sum_{i=0}^{p-1}\lambda^{i}(K^{p}+\mu K^{p})\subset K,$
and note that $r_{0}\in R$. Comparing coefficients of $S_{j}^{p^{n}}$ in (3.12), and using (3.13), we find that
$\alpha^{1-p+j}\in r_{0}^{-1}R,\hskip 21.68121pt0\leq j<p^{n},j\not\equiv-1\pmod{p}$.
Since $r_{0}^{-1}R$ is closed under multiplication by $K^{p}$, we also have $\alpha^{1+j}\in r_{0}^{-1}R$, hence $\alpha^{j}\in r_{0}^{-1}R$ for all $0<j<p$. Since $1\in r_{0}^{-1}R$, and $r_{0}^{-1}R$ is closed under multiplication by $\lambda$, it follows that $K^{p}[\alpha,\lambda]\subset r_{0}^{-1}R$. We claim that $\alpha\in K^{p}(\lambda)$. For if not, then – because $K$ has degree of imperfection $2$ – it would follow that $K=K^{p}[\alpha,\lambda]\subset r_{0}^{-1}R$, hence $R=K$. But $\mu^{2}\notin R$ due to the $p$-independence of $\lambda,\mu$. (Here we use $p>2$.) Thus
$\alpha\in K^{p}(\lambda),$ (3.15)
as claimed.
We claim that $F$ is homogeneous of degree $p^{n-1}$. To prove this, we first extend scalars to $L_{1}:=K(\mu^{1/p})$. Taking $Z,Z_{j}$ (instead of $S,S_{j}$) to be the variables on $V_{1,\lambda}$, we have the map $h_{1}\colon U_{L_{1}}\rightarrow V_{1,\lambda}$ defined by the formula $(Y,(X_{i})_{i})\mapsto(Z,(Z_{j})_{j})$ where $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for 0<j<p10<j<p-1, and $Z_{0}:=X_{0}+\mu^{1/p}Y$. Because $\alpha\notin L_{1}^{p}=K^{p}(\mu)$, due to (3.9) and the equality $K^{p}(\lambda)\cap K^{p}(\mu)=K^{p}$, we may Lemma 3.3 to the map $h_{1}\circ g$, in conjunction with [Ros1, Prop. 6.4] to conclude that $(X_{0}+\mu^{1/p}Y)(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$ (by which we mean once one excludes the linear term involving $S$). Now we extend scalars to $L_{2}:=K((\mu/\lambda)^{1/p})$. Over $L_{2}$, we have the map $h_{2}\colon U_{L_{2}}\rightarrow V_{1,\lambda}$ defined by the formula $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for $j=0$ and 1<j<p11<j<p-1, and $Z_{1}:=X_{1}+(\mu/\lambda)^{1/p}Y$. (Here we again use $p>2$ to ensure that 1<p11<p-1.) Once more, we may apply Lemma 3.3 to $h_{2}\circ g$ to conclude that $Z_{0}=X_{0}$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$. Since $X_{0}+\mu^{1/p}Y$ was as well, we deduce that $Y(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in those variables, hence $F$ is (see (3.5)). This proves the claim.
Now we are ready to obtain our contradiction and thereby complete the proof that $U$ is not unirational. Write $F=\sum_{j}\gamma_{j}S_{j}^{p^{n-1}}$. Comparing coefficients of $S_{j}^{p^{n}}$ in (3.12), and using (3.13), we find that $\mu(\gamma_{j}^{p}-\alpha^{1-p+j}c^{p})\in K^{p}[\alpha,\lambda]$. If $\gamma_{j}^{p}-\alpha^{1-p+j}c^{p}\neq 0$ for some $j$, then it follows that $\mu\in K^{p}[\alpha,\lambda]\subset K^{p}[\lambda]$ by (3.15), a contradiction. Thus we must have $\gamma_{j}^{p}-\alpha^{1-p+j}c^{p}=0$ for all $j$. Taking $j=0$, we must have $c=0$ because $\alpha\notin K^{p}$. Thus $\gamma_{j}=0$ for all $j$, hence $F=0$, in violation of (3.11) and (3.5).
Remark 3.6.
When $p=2$, the above argument does not work, and indeed, in this case $U$ is unirational, being a smooth quadric hypersurface (with a rational point). But if one instead takes the $K$-group
$U^{\prime}:=\{\mu Y^{4}-X_{1}+X_{0}^{2}+\lambda X_{1}^{2}=0\}\subset{\mathbf{G}}_{\rm{a}}^{3},$ (3.16)
then projection onto $Y$ defines a surjection $U^{\prime}\rightarrow{\mathbf{G}}_{\rm{a}}$ with unirational kernel, but $U^{\prime}$ is not unirational. We merely sketch the proof of this last claim. Using the notation of (3.5), one writes out the conditions for $Y,X_{j}$ to satisfy the defining equation (3.16) for $U^{\prime}$. One first shows that $\mathrm{deg}(F_{1})\leq 2^{n}$ as above, and then compares leading coefficients of $S_{j}^{2^{n+1}}$ and uses the $2$-independence of $\lambda,\mu$ and the fact that $\alpha\notin K^{2}$ to conclude that $c=0$. Then one shows that in fact $\mathrm{deg}(F_{1})\leq 2^{n-1}$. Because of the surjections $V_{n+1,\alpha}\twoheadrightarrow V_{n,\alpha}$ constructed in the proof of Lemma 3.3, one may assume that $n\geq 2$. Comparing coefficients of $S_{j}^{2^{n}}$, one then shows that $\alpha,\alpha^{-1}\in r_{0}^{-1}R^{\prime}$, where $R^{\prime}:=\mu K^{4}+K^{2}(\lambda)$ for some $0\neq r_{0}\in K^{2}(\lambda)$. One checks that this implies that in fact $\alpha\in K^{2}(\lambda)$. Then, analogously to the argument above, one passes to the extensions $L_{1}:=K(\mu^{1/2})$ and $L_{2}:=K((\mu/\lambda)^{1/4})$ and makes suitable changes of variables (same as in Example 3.5 for $L_{1}$, but in the case of $L_{2}$, the change is $(Y,X_{0},X_{1})\mapsto(Y,X_{0}+(\mu/\lambda)^{1/4}Y,X_{1}+(\mu/\lambda)^{1/2}Y^{2})$) to apply Lemma 3.3 and deduce that $F$ is homogeneous of degree $2^{n-2}$. Finally, one concludes as in the example above that in fact $F=0$, hence $Y=0$.
4 Ext-unirational groups
The examples of the preceding section demonstrate that unirationality is not inherited by extensions over fields of degree of imperfection >1>1. In light of this failure, it is natural to make the following definition.
Definition 4.1.
We say that a smooth $K$-group scheme $G$ is ext-unirational when it admits a filtration
$1=G_{0}\trianglelefteq G_{1}\trianglelefteq\dots\trianglelefteq G_{n}=G$
such that $G_{i+1}/G_{i}$ is unirational for all $0\leq i<n$.
Ext-unirational groups are necessarily affine and connected, because this is true of unirational groups. By Theorem 2.4, when $K$ has degree of imperfection $1$ (also, trivially, when $K$ is perfect), ext-unirationality is the same as unirationality, but over fields of larger degree of imperfection, ext-unirationality is a strictly weaker condition. Note also that ext-unirationality (like unirationality) is inherited by quotients and (unlike unirationality in general) by extensions. As usual, the notion of ext-unirationality is only interesting over imperfect fields, as over perfect fields every connected linear algebraic group is unirational.
Ext-unirationality may be rephrased more canonically as follows. Given a smooth finite type $K$-group scheme $G$, let $G_{\mathrm{uni}}\subset G$ denote its maximal unirational $K$-subgroup scheme. Then $G_{\mathrm{uni}}$ is a normal subgroup of $G$ [Ros3, Cor. 7.11]. Now define a sequence of $K$-group quotients $G_{\mathrm{uni}}^{[n]}$ of $G$ recursively by the formulas
$G_{\operatorname{uni}}^{[0]}:=G,\hskip 21.68121ptG_{\operatorname{uni}}^{[n+1]}:=G_{\operatorname{uni}}^{[n]}/(G_{\operatorname{uni}}^{[n]})_{\operatorname{uni}},n\geq 0.$
Then $G$ is ext-unirational if and only if $G_{\operatorname{uni}}^{[n]}=1$ for some (equivalently, all) $n\geq 0$. This also shows that we may replace the condition in Definition 4.1 that each $G_{i}$ be normal in $G_{i+1}$ by the condition that it is normal in $G$ without altering the definition. Furthermore, the construction of the $G_{\operatorname{uni}}^{[n]}$ commutes with separable extension of the ground field [Ros3, Cor. 7.10], and as a consequence, ext-unirationality is insensitive to passage to separable extensions: If $L/K$ is a separable field extension, then $G$ is ext-unirational over $K$ if and only if it is so over $L$.
Lemma 4.2.
If $G$ is a connected $K$-group scheme, then there is a surjection $f\colon G\twoheadrightarrow A$ such that $A$ is an abelian variety and $\ker(f)$ is affine.
Proof.
If $\mathrm{char}(K)=0$, then we are done by Chevalley’s Theorem. If $\mathrm{char}(K)>0$, then $G/I$ is smooth for some infinitesimal $K$-subgroup scheme $I\trianglelefteq G$ [SGA3, VIIA, Prop. 8.3], so we may assume that $G$ is smooth. By Chevalley again, $G$ admits a map as in the lemma over $K_{\operatorname{perf}}$, the perfect closure of $K$. Thus there is such a quotient over $G_{K^{1p^{n}}}$ for some $n>0$. We may identify $K^{1/p^{n}}$ with $K$ via the $p^{n}$-power map, and then $G_{K^{1/p^{n}}}$ becomes identified with $G^{(p^{n})}$. Thus $G^{(p^{n})}$ admits such a map. But because $G$ is smooth, $G^{(p^{n})}$ is the quotient of $G$ by the $n$th order relative Frobenius map over $K$, so $G$ also admits an abelian variety quotient with affine kernel. ∎
For a smooth connected $K$-group scheme $G$, we denote its derived group by $\mathscr{D}G$ and its abelianization $G/\mathscr{D}G$ by $G^{\operatorname{\rm{ab}}}$. Next we verify that ext-unirationality reduces to the $p$-torsion commutative setting.
Proposition 4.3.
A smooth connected $K$-group scheme $G$ is ext-unirational if and only if its maximal commutative quotient $G^{\operatorname{\rm{ab}}}$ is. If $G$ is affine and $\mathrm{char}(K)=p>0$, then this is further equivalent to ext-unirationality of its maximal $p$-torsion commutative quotient $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}$.
Proof.
When $\mathrm{char}(K)=0$, ext-unirationality is equivalent to affineness, so in this case it suffices to note that affineness of $G^{\operatorname{\rm{ab}}}$ implies the same for $G$, a consequence of Lemma 4.2 because abelian varieties are commutative. Since ext-unirationality implies affineness, this also reduces the assertion for smooth connected $G$ to the affine setting in characteristic $p$. Thus we may assume that $G$ is affine and that $\mathrm{char}(K)=p>0$, and we wish to show that ext-unirationality of $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}$ implies the same for $G$.
We may assume that $K$ is separably closed. The proof is by dimension induction. For any quotient $\overline{G}$ of $G$, the map $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}\rightarrow\overline{G}^{\operatorname{\rm{ab}}}/[p]\overline{G}^{\operatorname{\rm{ab}}}$ is surjective, so if $\mathrm{dim}(\overline{G})<\mathrm{dim}(G)$, then we may assume that $\overline{G}$ is ext-unirational. First suppose that $G=U$ is commutative and unipotent. If $U$ is $p$-torsion, then the proposition is immediate, so assume it is not, and let $n\geq 1$ be such that $p^{n+1}]U=0$ but $p^{n}]U\neq 0$. Let $\overline{U}:=U/[p^{n}]U$. Then $0<\mathrm{dim}(\overline{U})<\mathrm{dim}(U)$, and in particular, $\overline{U}$ is ext-unirational. The (nonzero) map $p]\colon U\rightarrow U$ factors through a nonzero map $p]\colon\overline{U}\rightarrow U$ whose image is ext-unirational. Since the image is nonzero, the cokernel is also ext-unirational, hence so is $U$. This completes the proof when $U$ is commutative unipotent.
Now suppose that $G=U$ is non-commutative unipotent, so the commutator map $c\colon U^{2}\rightarrow U$ is nonconstant. Let $\overline{U}$ be a maximal dimensional quotient of $U$ such that $c$ factors through a map $\overline{c}\colon\overline{U}\times U\rightarrow U$. Because $c$ is nonconstant, $\overline{U}\neq 1$. On the other hand, because $U$ is nilpotent, it admits a nontrivial smooth connected central $K$-subgroup $U^{\prime}$, and the commutator map factors through a map $U/U^{\prime}\times U\rightarrow U$, hence $\mathrm{dim}(\overline{U})<\mathrm{dim}(U)$. In particular, $\overline{U}$ is ext-unirational, so $\overline{U}_{\operatorname{uni}}\neq 1$. If the map $\overline{U}_{\operatorname{uni}}\rightarrow U$ defined by the formula $u\mapsto c(u_{1}u,u_{2})c(u_{1},u_{2})^{-1}$ is constant for all $u_{1}\in\overline{U}(K)$, $u_{2}\in U(K)$, then – because $K$ is separably closed and $\overline{U},U$ are smooth – it would follow that $c$ factors through a map $\overline{U}/\overline{U}_{\operatorname{uni}}\times U\rightarrow U$, in violation of the minimality of $\mathrm{dim}(\overline{U})$. Thus one can find $u_{1},u_{2}$ such that this map is nonconstant, so we have a nonconstant map from a positive-dimensional unirational $K$-scheme into $U$. It follows that $U_{\operatorname{uni}}\neq 1$, hence $U/U_{\operatorname{uni}}$ is ext-unirational, hence so is $U$. This completes the proof of the proposition for unipotent $G$.
Next we treat smooth connected affine $G$, where we assume that $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}$ is ext-unirational, and we wish to prove the same for $G$. Let $G_{t}\subset G$ denote the $K$-subgroup generated by the $K$-tori of $G$. This is a normal $K$-subgroup, and $U:=G/G_{t}$ is unipotent [CGP, Prop. A.2.11]. Furthermore, because tori are unirational, $G$ is ext-unirational if and only if $U$ is. The map $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}\rightarrow U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is surjective (in fact, it is an isomorphism) and ext-unirationality is inherited by quotients, so the general case follows from the unipotent one. ∎
When $K$ has degree of imperfection $1$, the above proposition has the following nice consequence.
Corollary 4.4.
Let $K$ be a field of degree of imperfection $1$. A smooth connected $K$-group scheme $G$ is unirational if and only if its abelianization $G^{\mathrm{ab}}$ is. If $G$ is also affine, then $G$ is unirational if and only if $G^{\operatorname{\rm{ab}}}/[p]G^{\operatorname{\rm{ab}}}$ is.
Proof.
Combine Proposition 4.3 and Theorem 2.4. ∎
We will discuss analogues of Corollary 4.4 in the next section.
Let $K$ be a field, and suppose given two smooth connected $K$-group schemes $G,H$. Let $\phi\colon G\dashrightarrow H$ be a rational map which is birational, with inverse $\psi\colon H\dashrightarrow G$. Let $U\subset G$ denote the open locus of definition of $\phi$. Denote by $f$ the composition $U\xrightarrow{\phi}H\xrightarrow{\pi_{H}}H/H_{\operatorname{uni}}$. We claim that $f$ factors through the open set $\pi_{G}(U)\subset G/G_{\operatorname{uni}}$, where $\pi_{G}\colon G\rightarrow G/G_{{}^{\prime}\operatorname{uni}}$ is the (flat) quotient map. For this assertion, we may assume that $K$ is separably closed, as the formation of the maximal unirational subgroup scheme is insensitive to separable extension. Let $a\colon U\times G_{\operatorname{uni}}\rightarrow G$ denote the action map $(u,x)\mapsto ux$, and let $V:=a^{-1}(U)$. By descent, $f$ factors through $\pi_{G}(U)$ if and only if the two maps $V\rightarrow H/H_{\operatorname{uni}}$, $(u,g)\mapsto\phi(u)$, $(u,g)\mapsto\phi(ug)$ coincide. We may verify this upon restriction to $V_{u_{0}}:=V\cap(u_{0}\times G_{\operatorname{uni}})$ for each $u_{0}\in U(K)$. Note that $V_{u_{0}}$ is open in $G_{\operatorname{uni}}$, hence unirational. We must show that the composition $V_{u_{0}}\rightarrow H\rightarrow H/H_{\operatorname{uni}}$ of $\pi_{H}\circ\phi$ is constant (where $\pi_{H}\colon H\rightarrow H/H_{\operatorname{uni}}$ is the quotient map), and this follows from the fact that the image of $\phi|V_{u_{0}}$ is a unirational subscheme of $H$. This proves that $f$ factors through a map $G/G_{\operatorname{uni}}\rightarrow H/H_{\operatorname{uni}}$. We may similarly factor the inverse map $\psi$ to get a map $H/H_{\operatorname{uni}}\rightarrow G/G_{\operatorname{uni}}$, and the compositions in both directions are the identity. We thereby deduce that $G_{\operatorname{uni}}$ is nontrivial if and only if $H_{\operatorname{uni}}$ is, and that $G/G_{\operatorname{uni}}$ is birationally equivalent to $H/H_{\operatorname{uni}}$. By dimension induction, therefore, $G$ is ext-unirational if and only if $H$ is. We have proved the following proposition.
Proposition 4.5.
If $G,H$ are smooth connected $K$-group schemes that are birationally equivalent, then $G$ is ext-unirational if and only if $H$ is.
If $T\subset G$ is a torus in a connected linear algebraic $K$-group $G$, then $G$ is unirational if and only if $Z_{G}(T)$ is [Ros3, Prop. 7.12]. Applying this with $T$ a maximal torus, and using Lemma 2.3, one sees that unirationality of $G$ is equivalent to that of the unipotent group $Z_{G}(T)/T$, whence of its maximal wound quotient. The same statement holds true for ext-unirationality, thanks to the following analogous result.
Proposition 4.6.
For a connected linear algebraic $K$-group $G$, the following are equivalent:
  • (i)
    $G$ is ext-unirational.
  • (ii)
    $Z_{G}(T)$ is ext-unirational for every $K$-torus $T\subset G$.
  • (iii)
    $Z_{G}(T)$ is ext-unirational for some $K$-torus $T\subset G$.
Proof.
We may assume that $K$ is separably closed. Let $T\subset G$ be a torus. For a generic cocharacter $\lambda\colon{\mathbf{G}}_{\rm{m}}\rightarrow T$ (lying in the complement of the union of finitely many hyperplanes in the cocharacter lattice), one has $Z_{G}(T)=Z_{G}(\lambda)$. Then there are (split) unipotent $K$-subgroups $U^{+},U^{-}\subset G$ such that the map
$U^{-}\times Z_{G}(T)\times U^{+}\rightarrow G,\hskip 21.68121pt(x,y,z)\mapsto xyz$
is an open embedding [CGP, Lem. 2.1.5,Props. 2.1.8(2)(3),2.1.10]. By Proposition 4.5, $G$ is ext-unirational if and only if $U^{-}\times Z_{G}(T)\times U^{+}$ is, which in turn holds if and only if $Z_{G}(T)$ is. ∎
5 Deducing unirationality from quotients
Corollary 4.4 says that, when $K$ has degree of imperfection $1$, unirationality of a group may be tested upon passage to a certain quotient of that group. In this section we will prove an analogue of this result for unipotent groups $U$ over fields of higher degree of imperfection. In particular, we will see that unirationality of $U^{\operatorname{\rm{ab}}}$ implies unirationality of $U$ when $K$ has degree of imperfection $\leq 2$. We will also show that this result is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational but $U$ is not (though $U$ must be ext-unirational by Proposition 4.3).
We first verify an analogous result for permawoundness in place of unirationality. The proof below makes reference to the notion of semiwoundness. Recall that a unipotent $K$-group scheme $U$ is called semiwound when it contains no copies of ${\mathbf{G}}_{\rm{a}}$ over $K$. This notion exhibits similar properties to the closely related notion of woundness, which is merely semiwoundness plus smoothness and connectedness; see [Ros2, Appendix A] for the basic properties of semiwound groups.
Proposition 5.1.
A smooth unipotent $K$-group scheme $U$ is permawound if and only if $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is.
Proof.
All smooth unipotent groups are permawound over perfect fields [Ros2, Prop. 5.2(i)], so we may assume that $K$ is imperfect. The only if direction follows from the fact that permawoundness is inherited by quotients, so we now concentrate on the converse. Let us first assume that $U$ is commutative (which, though we do not know it yet, must be the case a posteriori when $U$ is wound, as all wound permawound unipotent groups are commutative). The group $U$ is killed by $p^{n}$ for some $n\geq 0$, and we proceed by induction on $n$, the $n=0$ case being trivial. So suppose that $n>0$. The map $p^{n-1}]\colon U\rightarrow U$ descends to a surjective map $U/[p]U\twoheadrightarrow[p^{n-1}]U$, hence $p^{n-1}]U$ is permawound. If we let $\overline{U}:=U/[p^{n-1}]U$, then $U/[p]U\twoheadrightarrow\overline{U}/[p]\overline{U}$, so the latter group is permawound, hence, by induction, so is $\overline{U}$, because it is killed by $p^{n-1}$. Thus $U$ an extension of the two permawound groups $\overline{U}$ and $p^{n-1}]U$, hence is itself permawound by Proposition 2.1.
Now let $U$ be a semiwound unipotent group such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound. We will show that $U$ is commutative, hence permawound by the already-treated commutative case. First we note that $U$ must be connected, as otherwise $U$ would admit a nontrivial étale unipotent quotient, hence a nontrivial commutative $p$-torsion étale quotient, which would imply that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is disconnected, in violation of [Ros2, Prop. 6.2]. If $U=1$ the assertion is immediate, so assume that $U\neq 1$. Then $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}\subset U$. By dimension induction, the quotient $\overline{U}:=U/U^{\prime}$ is permawound, as $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ surjects onto $\overline{U}^{\operatorname{\rm{ab}}}/[p]\overline{U}^{\operatorname{\rm{ab}}}$. The commutator map $U^{2}\rightarrow U$ descends to a map $\overline{U}^{2}\rightarrow U$, which further descends to a map $\overline{U}_{w}^{2}\rightarrow U$, where $\overline{U}_{w}$ is the maximal wound quotient of $\overline{U}$. This map is constant by [Ros3, Cor. 10.3,Th. 1.9(ii)], hence $U$ is commutative, as claimed.
Finally, let $U$ be an arbitrary smooth unipotent $K$-group scheme such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound, and let $U_{s}\trianglelefteq U$ be the maximal split $K$-subgroup. Then $U/U_{s}$ is permawound by the semiwound case, hence so is $U$, by Proposition 2.1 and [Ros2, Prop. 5.3]. ∎
We will require the following result, of interest in its own right, that gives a finiteness criterion for permawoundness.
Proposition 5.2.
Let $K$ be a field of finite degree of imperfection. A smooth unipotent $K$-group scheme $U$ is permawound if and only if, for every commutative $p$-torsion wound unipotent $K$-group $V$, the group $\operatorname{Hom}_{K_{s}}(U,V)$ is finite.
Proof.
Thanks to Proposition 5.1, we may assume that $U$ is commutative and $p$-torsion. When $K$ is perfect, the only wound group is the trivial group, and every smooth unipotent $K$-group is permawound [Ros2, Prop. 5.2(i)], so we may assume that $K$ is imperfect. If $U_{s}\trianglelefteq U$ denotes the maximal split $K$-subgroup, then $U$ is permawound if and only if $U/U_{s}$ is, by Proposition 2.1 and [Ros2, Prop. 5.3]. Since $U_{s}$ is killed under any $K_{s}$-homomorphism from $U$ into a wound group, we may assume $U$ is semiwound (that is, contains no copy of ${\mathbf{G}}_{\rm{a}}$). The only if direction now follows from [Ros3, Prop. 10.1], plus the connectedness of $U$ (to ensure woundness of the maximal wound quotient) [Ros2, Prop. 6.2].
Now suppose that $K$ has finite nonzero degree of imperfection, and assume that $U$ is not permawound. We will show that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite, where (by abuse of notation) $\mathscr{V}$ is any $K$-form of the $K_{s}$-group which is denoted by $\mathscr{V}$ in [Ros2, Def. 7.3]. By [Ros2, Th. 1.4], we have for some $n>0$ (>0>0 because $U$ is not permawound) an exact sequence
$0\longrightarrow U\longrightarrow W\longrightarrow{\mathbf{G}}_{\rm{a}}^{n}\longrightarrow 0$
with $W$ wound, permawound, commutative, and $p$-torsion. We will show that $\operatorname{Ext}^{1}_{K_{s}}(W,\mathscr{V})$ is finite but $\operatorname{Ext}^{1}_{K_{s}}({\mathbf{G}}_{\rm{a}},\mathscr{V})$ is infinite, which will imply that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite as well. We may assume that $K=K_{s}$ [Ros3, Cor. 10.5].
First we prove the finiteness of $\operatorname{Ext}^{1}(W,\mathscr{V})$. We have a homomorphism $\phi\colon\operatorname{Ext}^{1}(W,\mathscr{V})\rightarrow\operatorname{Hom}(W,\mathscr{V})$ defined as follows: Given an extension $E$ in the former group, the Verschiebung endomorphism of $E$ descends to a map $W\rightarrow\mathscr{V}$ from the quotient $W$ to the subgroup $\mathscr{V}$. Then $\ker(\phi)$ consists of the extensions with trivial Verschiebung, and this is a trivial group by [Ros2, Cor. 8.3]. Thus the desired finiteness follows from the finiteness of $\operatorname{Hom}(W,\mathscr{V})$.
It remains to prove the infinitude of $\operatorname{Ext}^{1}({\mathbf{G}}_{\rm{a}},\mathscr{V})$. This group is a $K$-vector space via the action of $K$ on ${\mathbf{G}}_{\rm{a}}$, so it is equivalent to show that it is nonzero. By [CGP, Prop. B.1.13], we have an exact sequence
[…]
for a suitable $p$-polynomial $F$, where $d:=\mathrm{dim}(\mathscr{V})$. The sequence (5.1) does not split, because the wound $\mathscr{V}$ cannot be a quotient of ${\mathbf{G}}_{\rm{a}}^{d+1}$. Thus this sequence yields a nonzero element of $\operatorname{Ext}^{1}({\mathbf{G}}_{\rm{a}},\mathscr{V})$, as required. ∎
Proposition 5.3.
If $K$ is imperfect, then every extension of a unirational $K$-group scheme by a permawound unipotent $K$-group $U$ is unirational.
Proof.
We may assume that $K$ is separably closed. If $U$ is not semiwound, then we may use the fact that $\mathrm{H}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$ and dimension induction to conclude. We therefore assume that $U$ is semiwound. If $K$ has infinite degree of imperfection, then the assertion is immediate, thanks to [Ros2, Prop. 6.3]. Thus we may assume that $K$ has finite degree of imperfection. Because permawound groups are unirational [Ros3, Th. 1.9], we may use the rigidity property of permawound groups [Ros2, Th. 1.5] and an easy induction to reduce to showing that any torsor for the $K$-group $V$ over a unirational $K$-scheme $X$ is also unirational, where $V$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$. But this assertion follows from [Ros3, Prop. 8.6]. ∎
The proof of the following lemma uses the notion of the restricted moduli space of pointed morphisms $\mathscr{M}\mathit{or}((X,x),(G,1))^{+}$ from a $K$-scheme into a $K$-group scheme. This is the functor
$\{\mbox{geometrically reduced $K$-schemes}\}\rightarrow\{\mbox{groups}\}$
that sends a test scheme $T$ to the space of pointed $T$-morphisms $(X\times_{K}T,x_{T})\rightarrow(G\times_{K}T,1_{T})$. It turns out that it is represented by a smooth unipotent $K$-group scheme when $X$ is geometrically reduced of finite type and $G$ is wound unipotent, and in fact exhibits the somewhat stronger property that the corresponding scheme may be regarded as a subfunctor of the above morphism functor considered on the category of all $K$-schemes [Ros1, Th. 4.3].
Lemma 5.4.
Let $d,n,r\geq 0$ be integers.
  • (i)
    There is a constant $C=C(d,n,r)>0$ with the following property: Let $K$ be a field of characteristic $p$ and degree of imperfection $r$, and let $U\simeq\{F=0\}$ with $F$ a reduced $p$-polynomial over $K$ of degree $d$. Then for every $r$-tuple of divisors $D_{1},\dots,D_{r}\subset\mathbf{P}^{1}_{K}$ with $\sum_{i}\mathrm{deg}(D_{i})\leq n$, there are $\leq C$ multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$.
  • (ii)
    If $U$ is wound unipotent over a field $K$ of degree of imperfection $r$, then there is a constant $C=C(U,n)>0$ such that, for every $r$-tuple of divisors $D_{1},\dots,D_{r}\subset\mathbf{P}^{1}_{K}$ with $\sum_{i}\mathrm{deg}(D_{i})\leq n$, there are $\leq C$ multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$.
Proof.
Assertion (ii) follows from (i), since $U$ admits a filtration by commutative $p$-torsion wound unipotent groups by [CGP, Prop. B.3.2], and every such group is of the form $\{F=0\}$ for some reduced $p$-polynomial $F$ [CGP, Prop. B.1.13]. (We may assume that $K$ is infinite, since otherwise $r=0$.) So we concentrate on (i). We may assume that $K$ is separably closed. By [Ros3, Lem. 5.4], one may write $D_{i}$ as a finite disjoint union of $K$-schemes of the form $\mathrm{Spec}(K(\alpha_{ij}^{1/p^{n_{ij}}}))$ with $\alpha_{ij}\in K$, hence $\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{j}\operatorname{R}_{K(\alpha_{ij}^{1/p^{n_{ij}}})/K}({\mathbf{G}}_{\rm{m}})$, so we are reduced to the case in which $D_{i}=\mathrm{Spec}(K(\lambda_{i}^{1/p^{n_{i}}}))$ for all $i$.
[…]
We proceed by induction on $r$, the case $r=0$ being trivial. Suppose that $r>0$. By the ubiquity property of permawound groups [Ros2, Th. 1.4], there is an exact sequence
$0\longrightarrow U\longrightarrow W\longrightarrow{\mathbf{G}}_{\rm{a}}^{m}\longrightarrow 0$
with $W$ permawound. We require a slightly more precise result, namely, that $m$ may be bounded depending only on $r$ and $d$. (Note that $p=\mathrm{char}(K)$ is determined by $d$.) This follows from the proof of ubiquity. In particular, $\dim(W)$ may be bounded depending only on $r$ and $d$. By the rigidity property of permawound groups [Ros2, Th. 1.5], together with the fact that the number of terms in a filtration as in that result is bounded in terms of $\dim(W)$ and $p$, we may in fact assume that $U$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$ (even though technically the former is not wound, being nonsmooth). In the former case, $U$ is totally nonsmooth, so the only map from a geometrically reduced $K$-scheme into $U$ is the 0 map. So assume that $U=\mathscr{V}$, and we must prove the existence of a suitable $C=C(p,r,n_{1},\dots,n_{r})$.
We may assume that all $n_{i}$ are as small as possible – that is, if $n_{i}>0$, then $\lambda_{i}\notin K^{p}$. If $n_{i}=0$, then $X_{1}\simeq\mathbf{A}^{1}$, so any map $\prod X_{i}\rightarrow U$ which vanishes on $\infty\times\prod_{i>1}X_{i}$ must be constant because $U$ is wound. Thus we may assume that $\lambda_{1}\notin K^{p}$ and $n_{1}>0$. By [Ros3, Lem. 5.3], there is a totally nonsmooth (over $K$) $K$-subgroup $N\subset M:=\mathscr{M}\mathit{or}((\mathbf{P}^{1}\backslash\lambda_{1}^{1/p^{n_{1}}},\infty),(\mathscr{V},0))^{+}$ such that $(M_{K(\lambda_{1}^{1/p^{\infty}})})_{s}\subset N_{K(\lambda_{1}^{1/p^{\infty}})}$, where, for a unipotent $L$-group $V$, $V_{s}$ denotes the maximal split $L$-subgroup of $V$.
Now a map $\prod_{i}X_{i}\rightarrow U$ which vanishes whenever any coordinate is $\infty$ is the same thing as such a map $g\colon\prod_{i>1}X_{i}\rightarrow M$, and we will show that the number of such maps is finite and bounded depending only on the $n_{i}$, $r$, and $p$. Let $L:=K(\lambda_{1}^{1/p^{\infty}})$, a field of degree of imperfection r1r-1. Any map $g$ as above induces a map $\overline{g}\colon\prod_{i>1}X_{i}\rightarrow M_{L}/(M_{L})_{s}$ which vanishes whenever some coordinate is $\infty$. By induction, there are at most $C^{\prime}=C^{\prime}(d^{\prime},r-1,n_{2},\dots,n_{r})$ such maps, where $d^{\prime}$ is the degree of an equation describing $M_{L}/(M_{L})_{s}$. Further, we claim that if $\overline{g}=0$, then $g=0$. For if $\overline{g}=0$, then $g$ lands in $N$, since $N$ contains $(M_{L})_{s}$. Because $N$ is totally nonsmooth over $K$ and $\prod_{i>1}X_{i}$ is smooth, we must have $g=0$, as claimed.
It only remains to show that $d^{\prime}$ is bounded in terms only of $n_{1},r,p$. For this we argue as follows. Complete $\lambda_{1}$ to a $p$-basis $\mu_{1}:=\lambda_{1},\mu_{2},\dots,\mu_{r}$ of $K$, so in particular the $\mu_{i}$ are $p$-independent. Consider the subfield $F:=\mathbf{F}_{p}(\mu_{1},\dots,\mu_{r})\subset K$. Then $K/F$ is a separable extension, hence the formation of $M=\mathscr{M}\mathit{or}_{F}((\mathbf{P}^{1}\backslash\mu_{1}^{1/p^{n_{1}}},\infty),(\mathscr{V},0))^{+}$ commutes with scalar extension to $K$, where by abuse of notation, $\mathscr{V}$ is an $F$-descent of $\mathscr{V}/K$ (which exists because $\mathscr{V}$ may be defined via an equation using the $p$-basis $\mu_{1},\dots,\mu_{r}$ of $F$). Furthermore, if we let $E:=F(\lambda_{1}^{1/p^{\infty}})$, then $L/E$ is separable, hence $(M_{L})_{s}=(M_{E})_{s}\times_{E}L$. Thus we have reduced the calculation of $M_{L}/(M_{L})_{s}$ to the “universal” case over the single field $F$ with fixed $p$-basis $\mu_{1},\dots,\mu_{r}$. Thus the $d^{\prime}$ that works over this field also works in general, and has no dependence on the chosen $\lambda_{i}\in K$. The proof of the lemma is complete. ∎
Proposition 5.5.
Let $K$ be a field of degree of imperfection $r<\infty$, let $(X_{1},x_{1}),\dots,(X_{r},x_{r})$ be pointed unirational $K$-schemes with $x_{i}\in X_{i}(K)$, and let $U$ be a wound unipotent $K$-group. Then there are only finitely many maps $f\colon\prod_{i=1}^{r}X_{i}\rightarrow U$ such that $f$ vanishes whenever any one of the coordinates is set to $x_{i}$.
Proof.
We first treat the special case in which the $X_{i}$ are all open subschemes of $\mathbf{P}^{1}_{K}$, in which case we prove a somewhat stronger result – namely, that the number of maps $f$ as in the proposition is bounded in terms only of $U$ and the degrees of the complementary divisors $\mathbf{P}^{1}\backslash X_{i}$. As in the proof of Lemma 5.4 above, we may invoke [Ros1, Th. 6.7] to conclude that it is sufficient to show that, for any reduced divisors $D_{1},\dots,D_{r}$ on $\mathbf{P}^{1}$, the number of multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ is finite and bounded in terms only of $U$ and the degrees of the $D_{i}$, and this is the content of Lemma 5.4(ii).
Now consider the general case. For technical reasons, we shall find it more convenient to prove a slightly modified version of the proposition. We invoke certain notation from the beginning of [Ros3, §7]. Namely, we have associated to $f$ a map $\Delta_{f,x_{1},\dots,x_{r}}\colon\prod_{i=1}^{r}X_{i}\rightarrow U,$ which enjoys two key properties: (1) $\Delta_{f,x_{1},\dots,x_{r}}=1$ whenever any coordinate is set to $x_{i}$, and (2) if $f$ vanishes whenever any coordinate is set to $x_{i}$, then $\Delta_{f,x_{1},\dots,x_{r}}=f$. Further, the construction of $\Delta_{f,x_{1},\dots,x_{r}}$ is functorial in the obvious sense.
Choose dominant maps $g_{i}\colon Y_{i}\rightarrow X_{i}$ with $Y_{i}\subset\mathbf{P}^{m_{i}}$ nonempty open subschemes, and let $d_{i}$ denote the degree of the complementary divisor of $Y_{i}$. Let $C$ be such that there are at most $C$ maps as in the proposition whenever all of the $X_{i}$ are open subsets of $\mathbf{P}^{1}$ with complementary divisors whose degrees sum to $\sum_{i}d_{i}$. Given a set $S$ of >C>C maps $f\colon\prod_{i}X_{i}\rightarrow U$ such that $f$ vanishes whenever some coordinate is set to $x_{i}$, we will show that two of the $f\in S$ must coincide. For each $f$, let $g_{f}:=f\circ\prod_{i}g_{i}\colon\prod Y_{i}\rightarrow U$. For each $r$-tuple of points $y_{i}\in Y_{i}(K)$, consider the map $\Delta_{g_{f},y_{1},\dots,y_{r}}\colon\prod Y_{i}\rightarrow U$, which vanishes whenever some coordinate is set to $y_{i}$. For each $i$, let $L_{i}$ denote the space of lines on $\mathbf{P}^{m_{i}}$ through $y_{i}$. We may assume that $K$ is infinite (else $r=0$ and the proposition is trivial). Then for a Zariski dense set of points $\ell\in(\prod L_{i})(K)$, one has that $Y_{\ell}:=\ell\cap\prod Y_{i}$ is a product of open subschemes of $\mathbf{P}^{1}$ with complementary divisor on the $i$th coordinate of degree $d_{i}$. It follows from our choice of $C$ that two of the $\Delta_{g_{f},y_{1},\dots,y_{r}}$ must coincide when restricted to $Y_{\ell}$. Because the $Y_{\ell}$ are Zariski dense in $\prod Y_{i}$, we must therefore have that two of the $\Delta_{g_{f},y_{1},\dots,y_{r}}$ coincide.
Thus we have shown that for each $y\in\prod Y_{i}(K)$, there are two $f\in S$ such that $\Delta_{g_{f},y}$ coincide. It follows that there exist $f_{1},f_{2}\in S$ such that $\Delta_{g_{f_{1}},y}=\Delta_{g_{f_{2}},y}$ for a Zariski dense set of $y\in\prod Y_{i}(K)$. Because $\Delta_{g_{f},y_{1},\dots,y_{r}}=\Delta_{f,g(y_{1}),\dots,g(y_{r})}\circ g$, and $g$ is dominant, it follows that $\Delta_{f_{1},f_{1}(y_{1}),\dots,f_{1}(y_{r})}=\Delta_{f_{2},f_{2}(y_{1}),\dots,f_{2}(y_{r})}$ for a Zariski dense set of $\prod y_{i}\in\prod Y_{i}(K)$. Thus one has that $\Delta_{f_{1},x}=\Delta_{f_{2},x}$ for a Zariski dense set of $x\in\prod X_{i}(K)$, hence for all $x\in\prod X_{i}(K)$. In particular, this holds for $x=(x_{1},\dots,x_{r})$. Because, for all $\in S, ff$ vanishes whenever some coordinate is set to $x_{i}$, it follows that $f_{1}=f_{2}$. This proves the proposition. ∎
Proposition 5.6.
Let $K$ be a field of degree of imperfection $r<\infty$, let $(X_{1},x_{1}),\dots,(X_{r},x_{r})$ be pointed unirational $K$-schemes, and let $U$ be a wound unipotent $K$-group. Finally, let $f\colon\prod_{i=1}^{r}X_{i}\rightarrow U$ be such that $f$ vanishes whenever any one of the coordinates is set to $x_{i}$. Then $f$ generates a permawound $K$-subgroup of $U$.
Proof.
We may assume that $K$ is separably closed. We may additionally replace $U$ by the subgroup generated by $f$ and thereby assume that $f$ generates $U$. If $U$ is not permawound, then by Proposition 5.2 there is a wound unipotent $K$-group $V$ such that $\operatorname{Hom}(U,V)$ is infinite. But by postcomposing $f$ with these homomorphisms, we would thereby obtain a violation of Proposition 5.5. ∎
We are now prepared to prove the main result of this section. Let $G$ be a smooth connected $K$-group scheme. Recall that the lower central series $\mathscr{D}_{n}(G)$ of $G$ is defined by the formula $\mathscr{D}_{1}(G):=G$, $\mathscr{D}_{n+1}(G):=[G,\mathscr{D}_{n}(G)]$ for $n\geq 1$. This group must eventually stabilize, and we by define $\mathscr{D}_{\infty}(G)$ to be the stabilization.
Theorem 5.7.
Let $K$ be a field of degree of imperfection $r$, and let $U$ be a smooth connected unipotent $K$-group. Then $U$ is unirational if and only if $U/\mathscr{D}_{r}U$ is.
Proof.
Because $\mathscr{D}_{\infty}(U)=1$, we may assume that $r<\infty$. The only if direction is clear, so we prove the converse. If $U_{w}$ denotes the maximal wound quotient of $U$ (that is, the quotient of $U$ by its maximal split $K$-subgroup), then $U$ is unirational if and only if $U_{w}$ is, so we may assume that $U$ is wound. We may of course also assume that $U\neq 1$. Thus $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}$. Consider the $r$-fold commutator map $U^{r}\rightarrow U$, $(u_{1},\dots,u_{r})\mapsto[u_{1},[u_{2},\dots,u_{r}]\dots]$. This descends to a map $c\colon\overline{U}^{r}\rightarrow U$, where $\overline{U}:=U/U^{\prime}$. Because $\mathscr{D}_{r}U\twoheadrightarrow\overline{U}/\mathscr{D}_{r}(\overline{U}), U¯\overline{U}$ is unirational by dimension induction. If $c$ is constant, then $\mathscr{D}_{r}(U)=1$ and the theorem is trivially true. Otherwise, Proposition 5.6 implies that $c$ generates a nontrivial (normal) permawound $K$-subgroup $V$ of $U$. We then finish by combining dimension induction applied to $U/V$ and Proposition 5.3. ∎
Remark 5.8.
If $U$ is a wound unirational $K$-group, and $K$ has degree of imperfection $r$, then $\mathscr{D}_{r+1}(U)=1$ [Ros3, Th. 1.5]. (Note that the indexing on the lower central series used in that paper is unfortunately nonstandard, and in particular differs from that used here by $1$.) Therefore the quotient $U/\mathscr{D}_{r}(U)$ is in that sense not too far from $U$.
Of particular note in Theorem 5.7 is the case $r=2$, which says that if $K$ has degree of imperfection $2$, then $U$ is unirational if $U^{\operatorname{\rm{ab}}}$ is. We shall now give an example to show that this optimal, by constructing over every field of degree of imperfection $\geq 3$ a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not.
Example 5.9.
Let $K$ be a field of degree of imperfection $\geq 3$, and let $\lambda,\mu,\gamma\in K$ be $p$-independent. Recall that we have defined the following $K$-group
$V_{1,\gamma}:=\left\{-X_{p-1}+\sum_{i=0}^{p-1}\gamma^{i}X_{i}^{p}=0\right\}$
and $V_{1,\lambda}$, $V_{1,\mu}$ similarly. We also define the group
$W:=\left\{-Z_{p-1,p-1}+\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}Z_{i,j}^{p}=0\right\},$
where the variables in the $W$ equation are $Z_{i,j}$ for $0\leq i,j<p$. We have a bi-additive map
$b\colon V_{1,\lambda}\times V_{1,\mu}\rightarrow W$
defined by $(X_{i})_{i}\times(Y_{j})_{j}\mapsto(Z_{i,j}:=X_{i}Y_{j})_{i,j}$. That this does indeed land in $W$ is seen by the following calculation:
$Z_{p-1,p-1}=X_{p-1}Y_{p-1}=\left(\sum_{i=0}^{p-1}\lambda^{i}X_{i}^{p}\right)\left(\sum_{j=0}^{p-1}\mu^{j}Y_{j}^{p}\right)=\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}(X_{i}Y_{j})^{p}=\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}Z_{i,j}^{p}.$
We claim that $b$ generates $W$. That is, $b$ does not land in any proper $K$-subgroup scheme of $W$, or equivalently, for some $n>0$ the map $b^{n}\colon(V_{1,\lambda}\times V_{1,\mu})^{n}\rightarrow W$ obtained by adding $b$ over the $n$ coordinates is surjective. In fact, since all groups and maps live over the subfield $\mathbf{F}_{p}(\lambda,\mu)$, it suffices to verify this over this subfield, where it is [Ros3, Prop. 9.9]. (The multi-additive map given in the proof of that proposition is in this setting exactly $b$.)
Let
$G:=V_{1,\lambda}\times V_{1,\mu}\times V_{1,\gamma},$
and let $X:=W\times G$. For any bi-additive map $h\colon G\times G\rightarrow W$, we obtain a group structure $X_{h}$ on $X$ via $(w_{1},g_{1})\cdot(w_{2},g_{2}):=(w_{1}+w_{2}+h(g_{1},g_{2}),g_{1}+g_{2})$. The projection map $X_{h}\rightarrow G$ is then a homomorphism with kernel (isomorphic to) $W$, and this extension of $G$ by $W$ is central. Furthermore, one readily verifies that the commutator of $(0,g_{1}),(0,g_{2})$ is $h(g_{1},g_{2})-h(g_{2},g_{1})\in W$. We apply these observations to the bi-additive map $h_{0}\colon G^{2}=(V_{1,\lambda}\times V_{1,\mu}\times V_{1,\gamma})^{2}\rightarrow W$ defined by $h((v_{1},v_{2},v_{3}),(v_{1}^{\prime},v_{2}^{\prime},v_{3}^{\prime})):=b(v_{1},v_{2}^{\prime})$. Let $U_{1}:=X_{h_{0}}$. If we write $g_{i}:=(x_{i},y_{i},z_{i})$, then $(0,g_{1}),(0,g_{2})]=b(x_{1},y_{2})-b(x_{2},y_{1})$. In particular, setting $x_{2}=0$ and using the fact that $b$ generates $W$, we see that $U_{1}^{\operatorname{\rm{ab}}}=G$.
Now we define another extension $U_{2}$ of $G$ by $W$ as follows. By definition, the group $W$ sits in an exact sequence
[…]
where $F((Z_{i,j})_{ij}):=-Z_{p-1,p-1}+\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}Z_{i,j}^{p}.$ Associated to (5.2) we obtain a connecting map $\delta\colon\operatorname{Hom}(G,{\mathbf{G}}_{\rm{a}})\rightarrow\operatorname{Ext}^{1}(G,W)$. Let $\chi\colon G\rightarrow{\mathbf{G}}_{\rm{a}}$ denote the homomorphism $G\rightarrow V_{1,\gamma}\xrightarrow{X_{p-1}}{\mathbf{G}}_{\rm{a}}$, where the first map is projection. Then we define $U_{2}:=\delta(\chi)$. We claim that $U_{2}$ is not unirational. Let us grant this claim for the moment, and use it to construct our example of a unipotent $K$-group $U$ such that $U$ is not unirational but $U^{\operatorname{\rm{ab}}}$ is.
We define $U$ to be the Baer sum (which makes sense – and defines an abelian group structure – on the set of isomorphism classes of central extensions) of the two extensions $U_{1}$ and $U_{2}$ of $G$ by $W$. That is, $U:=\operatorname{coker}(\Delta\colon W\rightarrow U_{1}\times_{G}U_{2})$, where $\Delta$ is the antidiagonal map $w\mapsto(w,-w)$. We first check that $U$ is not unirational. Indeed, the forgetful map $\operatorname{Ext}^{1}(G,W)\rightarrow{\rm{H}}^{1}(G,W)$ which recalls only the structure of the extension as a $W$-torsor over $G$ is a homomorphism. Since $U_{1}$ is by definition trivial as a torsor, it follows that $U\simeq U_{2}$ as a $W$-torsor over $G$. In particular, they are isomorphic as $K$-schemes. Since the latter is not unirational, therefore, neither is the former. Now a straightforward calculation (most easily carried out by regarding the derived group as an fppf sheaf rather than as an algebraic group) using the fact that $W\subset\mathscr{D}U_{1}$ shows that $W\subset\mathscr{D}U$, hence $U^{\operatorname{\rm{ab}}}=G$ is unirational.
It remains to prove that $U_{2}$ is not unirational. The argument is more or less the same as in Example 3.4. As in that example, it is sufficient to show that there is no nonzero homomorphism (with $n>0$ and $\alpha\in K-K^{p}$) $f\colon V_{n,\alpha}\rightarrow V_{1,\gamma}$ such that $X_{p-1}\circ f$ lifts through the map $F$ in (5.2) to a homomorphism $V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}^{p^{2}}$. (See (3.1) for the definition of $V_{n,\alpha}$.) Suppose to the contrary that there were. Then by Lemma 3.3, we would have $\alpha\in K^{p}(\gamma)$, and using $S,S_{k}$ to denote the variables on $V_{n,\alpha}$, we may write
$X_{p-1}\circ f=cS,\hskip 14.45377pt0\neq c\in K.$ (5.3)
Then we have an equation
$cS=-Z_{p-1,p-1}+\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}Z_{i,j}^{p},$ (5.4)
with each $Z_{i,j}$ a $p$-polynomial in $S$ and the $S_{k}$ of degree $\leq 1$ in $S$. We claim that $\mathrm{deg}(Z_{i,j})\leq p^{n-1}$. Indeed, using the equation (3.1) to eliminate $S^{p}$, (5.4) becomes an identity of polynomials. If some $Z_{i,j}$ has degree $p^{n-1}$, then looking at the leading term in some $S_{k}$ of (5.4) and using the $p$-independence of $\lambda,\mu$ would yield a contradiction. Thus we may write
$Z_{i,j}=d_{i,j}S+\sum_{\begin{subarray}{c}0\leq k<p^{n}\\ k\not\equiv-1\pmod{p}\end{subarray}}d_{i,j,k}S_{k}^{p^{n-1}}+G_{i,j},$ (5.5)
where $d_{i,j},d_{i,j,k}\in K$ and $G_{i,j}$ is a $p$-polynomial over $K$ in the $S_{k}$ (not involving $S$) of degree $p^{n-1}$. Then using (3.1) to eliminate $S^{p}$, and substituting (5.5) into (5.4) and comparing coefficients of $S_{0}^{p^{n}}$ yields
$\alpha^{1-p}\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}d_{i,j}^{p}=\sum_{0\leq i,j<p}\lambda^{i}\mu^{j}d_{i,j,0}^{p}.$
If not all $d_{i,j}=0$, then it would follow that $\alpha\in K^{p}(\lambda,\mu)$. But we have that $\alpha\in K^{p}(\gamma)$, so it would then follow (because $\lambda,\mu,\gamma$ are $p$-independent) that $\alpha\in K^{p}$, which is false. Thus $d_{i,j}=0$ for all $i,j$. Combining (5.5) and (5.4) then yields that $c=0$, in violation of (5.3). This completes the proof that $U_{2}$ is not unirational.

### v3

@@
 1 Introduction
+Recall that a finite type $K$-scheme $X$ is said to be unirational when there is a dominant rational map $\mathbf{A}^{n}\rightarrow X$ for some $n>0$.
+By dominant, we mean in the sense that the schematic closure is all of $X$, so in particular $X$ must be geometrically integral.
+Unirational varieties are of interest for a variety of reasons, one of which is that they admit many (a Zariski dense set of) rational points, at least when $K$ is infinite.
+Unirational algebraic groups, in particular, have received attention from various authors; see, for instance, [BLR, §10.3] (on the relationship between unirationality and the existence of Néron models for algebraic groups), [CGP, §11.3] (largely on the relationship between unirationality and various arithmetic properties of algebraic groups), [Ach2] (regarding geometric properties of unirational groups), [Ros1, §6] (about the Picard groups of unirational groups), and [Ros4] (concerned primarily with various rigidity properties exhibited by unirational groups, in some sense analogous to the well-known rigidity properties of abelian varieties).
+Over perfect fields, every connected linear algebraic groups is unirational [Bor, Ch.
+V,Th.
+18.2(ii)], but this is false over every imperfect field [CGP, Ex.
+11.3.1], and so unirationality becomes a nontrivial condition on (connected) linear algebraic groups in the imperfect setting.
+If $G$ is a connected linear algebraic group, and $T\subset G$ is a maximal $K$-torus, then $G$ is unirational precisely when the unipotent group $Z_{G}(T)/T$ is, where $Z_{G}(T)$ denotes the centralizer of $T$.
+(Combine [Ros4, Prop.
+7.12] and Lemma 2.3 below.) For this reason (and others), unirationality questions for algebraic groups often reduce to the unipotent setting, and thus the theory of unipotent groups plays a central role in such problems.
 In [Ach1, Question 4.9], Achet posed the following question: Is a commutative extension of two unirational algebraic $K$-groups necessarily also unirational?
-His motivation lies in the study of the maximal unirational $K$-subgroup $G_{\mathrm{uni}}$ of an algebraic group, as an affirmative answer implies that, for commutative $G$, the group $G/G_{\mathrm{uni}}$ admits no nontrivial unirational subgroup.
-The question trivially admits an affirmative answer when $K$ is perfect, as in this case all connected linear algebraic $K$-groups are unirational [Bor, Ch.
-V,Th.
-18.2(ii)].
+His motivation lies in the study of the maximal unirational $K$-subgroup $G_{\mathrm{uni}}$ of an algebraic group.
+(This subgroup of $G$ may also be described as the group generated by all maps from unirational varieties into $G$ which pass through the identity.) Achet posed his question because an affirmative answer implies that, for commutative $G$, the group $G/G_{\mathrm{uni}}$ admits no nontrivial unirational subgroup.
+The question trivially admits an affirmative answer when $K$ is perfect, since all connected linear algebraic groups are unirational in this case, but over imperfect fields it becomes highly nontrivial.
 In this paper, we answer Achet’s question, and more generally undertake a study of extensions of unirational algebraic groups.
 In §2, we prove that every extension of unirational groups is unirational over fields of degree of imperfection $1$ (Theorem 2.4).
-However, this fails over every field of degree of imperfection >1>1.
+This makes crucial use of the concept of permawound unipotent groups introduced in [Ros3].
+We will remind the reader of the definition of permawoundness in §2, but their utility arises not from their definition, but from the fact that they simultaneously exhibit two important properties: One the one hand, they enjoy a certain universality or ubiquity – every commutative $p$-torsion wound unipotent group may be embedded into a permawound one – so that many questions about arbitrary unipotent groups reduce to the permawound case.
+On the other hand, they exhibit an extraordinary rigidity: After passage to a suitable finite separable extension, every permawound group admits a filtration with successive quotients all isomorphic to one of two particular groups, very much analogously to the manner in which split unipotent groups always (by definition) admit a filtration with ${\mathbf{G}}_{\rm{a}}$ quotients [Ros3, Th.
+1.4,1.5].
+One sign that they are relevant for our investigations here is that such groups are always unirational [Ros4, Th.
+1.9(i)].
+In fact, we shall show that a smooth unipotent group over a field of degree of imperfection $1$ is unirational precisely when it is permawound (Proposition 2.2), from which the stability of unirationality under extensions will follow easily.
+However, this stability fails over every field of degree of imperfection >1>1.
 In fact, in §3 we construct over every such field a commutative $p$-torsion extension of unirational wound unipotent groups which is not unirational (Example 3.4).
@@
 Finally, in §5, we continue with the theme of deducing unirationality properties from quotients, but for unirationality itself rather than ext-unirationality.
-More precisely, we prove that over a field $K$ of degree of imperfection $r$, a smooth connected unipotent $K$-group $U$ is unirational if and only if the quotient $U/\mathscr{D}_{r}U$ of $U$ by the $r$th term in its lower central series is (Theorem 5.7).
-Of particular note is the case $r\leq 2$, which tells us that, when $K$ has degree of imperfection $\leq 2$, unirationality of $U$ is equivalent to that of its maximal commutative quotient $U^{\operatorname{\rm{ab}}}$.
-We show that this is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent $K$-group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not (Example 5.9).
+More precisely, we prove that over a field $K$ of degree of imperfection $r>1$, a smooth connected unipotent $K$-group $U$ is unirational if and only if the quotient $U/\mathscr{D}_{r}U$ of $U$ by the $r$th term in its lower central series is (Theorem 5.7).
+Of particular note is the case $r\leq 2$, which – in conjunction with Corollary 4.5 and the unirationality of all connected linear algebraic groups when $r=0$ – tells us that, when $K$ has degree of imperfection $\leq 2$, unirationality of $U$ is equivalent to that of its maximal commutative quotient $U^{\operatorname{\rm{ab}}}$.
+We show that this is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent $K$-group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational.
+(Example 5.9).
 Notation and Conventions
@@
 This quantity is always either infinite or a nonnegative integer.
-The degree of imperfection of $K$ also equals $\mathrm{dim}_{K}(\Omega^{1}_{K/\mathbf{F}_{p}})$ as well as the maximal size of a $p$-independent subset of $K$ [Mat, Th.
+The degree of imperfection of $K$ also equals $\mathrm{dim}_{K}(\Omega^{1}_{K/\mathbf{F}_{p}})$, as well as (tautologically, from the first definition) the size of any $p$-basis of $K$ [Mat, Th.
 26.5].
+Acknowledgments
+I thank Anis Zidani for pointing out an error in an earlier version of Theorem 5.7.
 2 Degree of imperfection 11
-In this section we will prove that an arbitrary (not necessarily commutative) extension of unirational algebraic $K$-groups must be unirational when $K$ has degree of imperfection $1$.
-We begin by tying up a loose end concerning a basic permanence property of permawound unipotent groups.
-Recall that a smooth unipotent $K$-group scheme $U$ is called permawound when, for any exact sequence of finite type $K$-group schemes
+In this section, we will prove that an arbitrary (not necessarily commutative) extension of unirational algebraic $K$-groups must be unirational when $K$ has degree of imperfection $1$.
+We begin by tying up a loose end concerning a basic permanence property of permawound unipotent groups, which were introduced in [Ros3].
+Recall [Ros3, Def.
+1.2] that a smooth unipotent $K$-group scheme $U$ is called permawound when, for any exact sequence of finite type $K$-group schemes
 $U\longrightarrow E\longrightarrow{\mathbf{G}}_{\rm{a}}\longrightarrow 1,$
-$E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$.
+$E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$; that is, there is an injective (or equivalently, there is a nonzero) $K$-homomorphism ${\mathbf{G}}_{\rm{a}}\rightarrow E$.
+As we shall show below (Proposition 2.2), it is an easy consequence of the results of [Ros3] and [Ros4] that permawoundness is equivalent to unirationality for unipotent groups over fields of degree of imperfection $1$.
+At least for unipotent groups, then, we will deduce the permanence of unirationality under extensions from the corresponding property of permawoundness, to be proved presently.
+The case of general algebraic groups will then be obtained by reduction to the unipotent case, using the fact, a consequence of [Ros4, Prop.
+7.12] and Lemma 2.3 below, that $G$ is unirational if and only if the same holds for $Z_{G}(T)/T$, where $T\subset G$ is a maximal torus.
 Proposition 2.1.
@@
 Proof.
-If $K$ is perfect, then the assertion is trivially true by [Ros2, Prop.
+If $K$ is perfect, then the assertion is trivially true by [Ros3, Prop.
 5.2(i)], so assume that $K$ is imperfect.
-Then $U^{\prime}$ and $U^{\prime\prime}$ are connected [Ros2, Prop.
+Then $U^{\prime}$ and $U^{\prime\prime}$ are connected [Ros3, Prop.
 6.2], hence so is $U$.
@@
 Because $\overline{U}^{2}(K_{s})$ is Zariski dense in $\overline{U}^{2}$, it follows that $c$ is $\overline{U}_{s}$-invariant in both arguments, hence $c$ indeed descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$ which we by abuse of notation also denote by $c$.
-Because $\overline{U}_{w}$ is wound and permawound, it is commutative [Ros3, Cor.
-10.4], so $c$ is the constant map to $1\in U(K)$ by [Ros3, Lem.
+Because $\overline{U}_{w}$ is wound and permawound, it is commutative [Ros4, Cor.
+10.4], so $c$ is the constant map to $1\in U(K)$ by [Ros4, Lem.
 10.2].
 It follows that $U$ is commutative.
-Because commutative permawound groups are trivially weakly permawound [Ros2, Def.
-5.1], $U^{\prime}$ and $U^{\prime\prime}$ are weakly permawound, hence so is $U$ [Ros2, Prop.
+Because commutative permawound groups are trivially weakly permawound [Ros3, Def.
+5.1], $U^{\prime}$ and $U^{\prime\prime}$ are weakly permawound, hence so is $U$ [Ros3, Prop.
 5.6].
-It follows that $U$ is permawound [Ros2, Prop.
+It follows that $U$ is permawound [Ros3, Prop.
 6.9].
@@
 Proof.
-For the if direction, use [Ros3, Th.
+For the if direction, use [Ros4, Th.
 1.9(i)].
-For the converse, [Ros3, Prop.
+For the converse, [Ros4, Prop.
 9.6] yields the result when $U$ is wound.
@@
 $1\longrightarrow U_{s}\longrightarrow U\longrightarrow U_{w}\longrightarrow 1$
-with $U_{s}$ split and $U_{w}$ wound. Because $U$ is unirational, so is $U_{w}$, which is therefore permawound. Then [Ros2, Prop. 5.3] and Proposition 2.1 show that $U$ is permawound. ∎
+with $U_{s}$ split and $U_{w}$ wound. Because $U$ is unirational, so is $U_{w}$, which is therefore permawound. Then [Ros3, Prop. 5.3] and Proposition 2.1 show that $U$ is permawound. ∎
 We require the following simple lemma.
@@
 Proof.
-We may assume that $K$ is separably closed [Ros3, Th.
+We may assume that $K$ is separably closed [Ros4, Th.
 1.6].
@@
 ∎
+We come now to the main result of this section.
 Theorem 2.4.
@@
 2].
-The groups $Z_{G^{\prime\prime}}(T^{\prime\prime})$ and $Z_{G^{\prime}}(T^{\prime})$ are unirational [Ros3, Prop 7.12], hence so are the two groups on the end in (2.1).
+The groups $Z_{G^{\prime\prime}}(T^{\prime\prime})$ and $Z_{G^{\prime}}(T^{\prime})$ are unirational [Ros4, Prop 7.12], hence so are the two groups on the end in (2.1).
 They are also unipotent, the unipotence a consequence of the maximality of the respective tori combined with [Bor, Ch.
@@
 Now an extension of unirational unipotent groups over a field of degree of imperfection $1$ is still unirational by combining Propositions 2.2 and 2.1.
-Thus $Z_{G}(T)/T$ is unirational, hence so is $Z_{G}(T)$ (Lemma 2.3), hence so is $G$ [Ros3, Prop.
+Thus $Z_{G}(T)/T$ is unirational, hence so is $Z_{G}(T)$ (Lemma 2.3), hence so is $G$ [Ros4, Prop.
 7.12].
@@
 In this section we will show by example – over every field of degree of imperfection >1>1 – that unirationality is not inherited by extensions in general.
-Our examples will be of two main types: Our first example will be an extension of a wound unirational group by a wound unirational group, while our second will be an extension of ${\mathbf{G}}_{\rm{a}}$ by a wound unirational group.
+Our examples will be of two types: Our first example will be an extension of a wound unirational group by a wound unirational group, while our second will be an extension of ${\mathbf{G}}_{\rm{a}}$ by a wound unirational group.
 Of course, any extension of a unirational group by a split unipotent one is unirational, as ${\rm{H}}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$.
@@
 This also may be seen as folllows: $V_{n,\alpha}$ is defined over the field $\mathbf{F}_{p}(\alpha)$, so we may verify its unirationality over this field, which has degree of imperfection $1$.
-But over $\mathbf{F}_{p}(\alpha)$, $V_{n,\alpha}$ is permawound [Ros2, Props.
-6.4,6.9,3.5], hence unirational [Ros3, Th.
+But over $\mathbf{F}_{p}(\alpha)$, $V_{n,\alpha}$ is permawound [Ros3, Props.
+6.4,6.9,3.5], hence unirational [Ros4, Th.
 1.9(i)].
@@
 It remains to prove that this (and indeed, any such) homomorphism is an isomorphism.
-The principal part of the polynomial (3.1) defining $V$ is reduced and universal over $F$ (for instance, using [Ros2, Prop.
-3.5(ii)]), so $V$ is permawound over $F$ [Ros2, Props.
-6.4,6.9], hence unirational [Ros3, Th.
+The principal part of the polynomial (3.1) defining $V$ is reduced and universal over $F$ (for instance, using [Ros3, Prop.
+3.5(ii)]), so $V$ is permawound over $F$ [Ros3, Props.
+6.4,6.9], hence unirational [Ros4, Th.
 1.9(i)].
@@
 To prove that $g$ is injective, we note that dimension considerations imply that $\ker(g)$ is finite.
-Because $V/\ker(g)\simeq\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is wound unipotent, $\ker(g)$ is weakly permawound [Ros2, Prop.
-5.5], hence has strictly positive dimension if nontrivial (for instance, by [Ros2, Th.
-9.5]; recall that weak permawoundness is insensitive to algebraic separable extension of the ground field [Ros2, Prop.
+Because $V/\ker(g)\simeq\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is wound unipotent, $\ker(g)$ is weakly permawound [Ros3, Prop.
+5.5], hence has strictly positive dimension if nontrivial (for instance, by [Ros3, Th.
+9.5]; recall that weak permawoundness is insensitive to algebraic separable extension of the ground field [Ros3, Prop.
 6.7]).
@@
 2].
-(For a more detailed explanation of how this follows from that theorem, see the discussion before Theorem 2.2 in [Ros1, §2].) Because $U$ is wound, $D$ may be taken to be reduced by [Ros1, Lem.
+(For a more detailed explanation of how this follows from that theorem, see the discussion before Theorem 2.2 in [Ros2, §2].) Because $U$ is wound, $D$ may be taken to be reduced by [Ros2, Lem.
 2.3].
@@
 Hence $U$ is generated by maps from $\operatorname{R}_{K(x)/K}({\mathbf{G}}_{\rm{m}})$ as $x$ varies over closed points of $\mathbf{P}^{1}$.
-Because $K$ is separably closed, $K(x)/K$ is a (finite) purely inseparable extension, and it is a primitive extension (generated by a single element) by [Ros3, Lem.
+Because $K$ is separably closed, $K(x)/K$ is a (finite) purely inseparable extension, and it is a primitive extension (generated by a single element) by [Ros4, Lem.
 5.4].
@@
 $W:=\left\{\sum_{i=0}^{p-1}\alpha^{i}X_{i}^{p}=0\right\}\subset{\mathbf{G}}_{\rm{a}}^{p},$
-which is isomorphic to $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ [Ros2, Prop. 7.4].
-By [Ros2, Lem.
+which is isomorphic to $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ [Ros3, Prop. 7.4].
+By [Ros3, Lem.
 7.9], therefore, and an easy induction, any map $V_{n,\alpha}\rightarrow V_{1,\lambda}$ must factor through the map $V_{n,\alpha}\rightarrow V_{1,\alpha}$ obtained by iterating the maps $f_{m}$ above.
 The lemma therefore reduces to the case $n=1$, where we must show that the map in question is obtained by a linear change of variables, and that the coefficient of $S$ in $X$ is nonzero.
-(That $X$ may be written uniquely as a $p$-polynomial of degree $\leq 1$ in $S$ follows from [Ros1, Prop.
+(That $X$ may be written uniquely as a $p$-polynomial of degree $\leq 1$ in $S$ follows from [Ros2, Prop.
 6.4].) For this, we may extend scalars to $K_{s}$ and thereby assume that $K$ is separably closed.
-The lemma then follows from [Ros2, Lem.
+The lemma then follows from [Ros3, Lem.
 7.1,Prop.
@@
 $f\circ g=cS\mbox{ for some }0\neq c\in K.$ (3.6)
-Applying [Ros1, Prop. 6.4], there exist $p$-polynomials $X_{0},\dots,X_{p-1}\in K[S,(S_{j})_{j}]$ of degree $\leq 1$ in $S$ such that
+Applying [Ros2, Prop. 6.4], there exist $p$-polynomials $X_{0},\dots,X_{p-1}\in K[S,(S_{j})_{j}]$ of degree $\leq 1$ in $S$ such that
 $f\circ g=cS=-X_{p-1}+\sum_{i=0}^{p-1}\mu^{i}X_{i}^{p}$ (3.7)
 as homomorphisms $V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}$.
-Applying [Ros1, Prop.
+Applying [Ros2, Prop.
 6.4] once more, if we use the equation (3.1) to eliminate $S^{p}$ from the $X_{i}^{p}$ terms above, then this becomes an identity of polynomials in $S$ and the $S_{j}$.
@@
 26.6].
-Thus the non-unirationality of $U$ may be checked over $\mathbf{F}_{p}(\lambda,\mu)$ [Ros3, Th.
+Thus the non-unirationality of $U$ may be checked over $\mathbf{F}_{p}(\lambda,\mu)$ [Ros4, Th.
 1.6], hence we may assume that $\lambda,\mu$ form a $p$-basis for $K$.
@@
 and a homomorphism $g\colon V_{n,\alpha}\rightarrow U$ whose projection onto the $Y$ coordinate is nonzero.
-By [Ros1, Prop.
+By [Ros2, Prop.
 6.4], $g=(Y(S,(S_{j})_{j}),X_{0}(S,(S_{j})_{j}),\dots,X_{p-1}(S,(S_{j})_{j}))$, where
@@
 $c_{p-1}S+\alpha^{1-p}\left(\mu c^{p}+\sum_{i=0}^{p-1}\lambda^{i}c_{i}^{p}\right)\left[S-\sum_{\begin{subarray}{c}0\leq j<p^{n}\\ j\not\equiv-1\pmod{p}\end{subarray}}\alpha^{j}S_{j}^{p^{n}}\right]-F_{p-1}+\mu F^{p}+\sum_{i=0}^{p-1}\lambda^{i}F_{i}^{p}=0.$ (3.12)
-Because the above expression has degree <p<p in $S$, it is an identity of $p$-polynomials in the variables $S,S_{j}$ [Ros1, Prop. 6.4]. We claim that
+Because the above expression has degree <p<p in $S$, it is an identity of $p$-polynomials in the variables $S,S_{j}$ [Ros2, Prop. 6.4]. We claim that
 $\mathrm{deg}(F_{p-1})\leq p^{n-1}.$ (3.13)
@@
 Taking $Z,Z_{j}$ (instead of $S,S_{j}$) to be the variables on $V_{1,\lambda}$, we have the map $h_{1}\colon U_{L_{1}}\rightarrow V_{1,\lambda}$ defined by the formula $(Y,(X_{i})_{i})\mapsto(Z,(Z_{j})_{j})$ where $Z:=X_{p-1}$, $Z_{j}:=X_{j}$ for 0<j<p10<j<p-1, and $Z_{0}:=X_{0}+\mu^{1/p}Y$.
-Because $\alpha\notin L_{1}^{p}=K^{p}(\mu)$, due to (3.9) and the equality $K^{p}(\lambda)\cap K^{p}(\mu)=K^{p}$, we may Lemma 3.3 to the map $h_{1}\circ g$, in conjunction with [Ros1, Prop.
+Because $\alpha\notin L_{1}^{p}=K^{p}(\mu)$, due to (3.9) and the equality $K^{p}(\lambda)\cap K^{p}(\mu)=K^{p}$, we may Lemma 3.3 to the map $h_{1}\circ g$, in conjunction with [Ros2, Prop.
 6.4] to conclude that $(X_{0}+\mu^{1/p}Y)(S,(S_{j})_{j})$ is homogeneous of degree $p^{n-1}$ in the $S_{j}$ (by which we mean once one excludes the linear term involving $S$).
@@
 Given a smooth finite type $K$-group scheme $G$, let $G_{\mathrm{uni}}\subset G$ denote its maximal unirational $K$-subgroup scheme.
-Then $G_{\mathrm{uni}}$ is a normal subgroup of $G$ [Ros3, Cor.
+Then $G_{\mathrm{uni}}$ is a normal subgroup of $G$ [Ros4, Cor.
 7.11].
-Now define a sequence of $K$-group quotients $G_{\mathrm{uni}}^{[n]}$ of $G$ recursively by the formulas
-$G_{\operatorname{uni}}^{[0]}:=G,\hskip 21.68121ptG_{\operatorname{uni}}^{[n+1]}:=G_{\operatorname{uni}}^{[n]}/(G_{\operatorname{uni}}^{[n]})_{\operatorname{uni}},n\geq 0.$
-Then $G$ is ext-unirational if and only if $G_{\operatorname{uni}}^{[n]}=1$ for some (equivalently, all) $n\geq 0$.
+Now define a sequence of $K$-group quotients $G_{u}^{[n]}$ of $G$ recursively by the formulas
+$G_{u}^{[0]}:=G,\hskip 21.68121ptG_{u}^{[n+1]}:=G_{u}^{[n]}/(G_{u}^{[n]})_{\operatorname{uni}},n\geq 0.$
+Then $G$ is ext-unirational if and only if $G_{u}^{[n]}=1$ for some (equivalently, all sufficiently large) $n\geq 0$.
 This also shows that we may replace the condition in Definition 4.1 that each $G_{i}$ be normal in $G_{i+1}$ by the condition that it is normal in $G$ without altering the definition.
-Furthermore, the construction of the $G_{\operatorname{uni}}^{[n]}$ commutes with separable extension of the ground field [Ros3, Cor.
-7.10], and as a consequence, ext-unirationality is insensitive to passage to separable extensions: If $L/K$ is a separable field extension, then $G$ is ext-unirational over $K$ if and only if it is so over $L$.
+Furthermore, the construction of the $G_{u}^{[n]}$ commutes with separable extension of the ground field [Ros4, Cor.
+7.10], and as a consequence, ext-unirationality is insensitive to passage to separable extensions:
+Proposition 4.2.
+If $L/K$ is a separable field extension, then $G$ is ext-unirational over $K$ if and only if it is so over $L$.
+The following lemma will be useful for proving results beyond the affine setting.
 Lemma 4.3.
@@
 Combine Proposition 4.4 and Theorem 2.4. ∎
-We will discuss analogues of Corollary 4.4 in the next section.
+We will discuss analogues of Corollary 4.5 over fields of higher degree of imperfection in the next section.
 Let $K$ be a field, and suppose given two smooth connected $K$-group schemes $G,H$.
@@
 If $G,H$ are smooth connected $K$-group schemes that are birationally equivalent, then $G$ is ext-unirational if and only if $H$ is.
-If $T\subset G$ is a torus in a connected linear algebraic $K$-group $G$, then $G$ is unirational if and only if $Z_{G}(T)$ is [Ros3, Prop.
+If $T\subset G$ is a torus in a connected linear algebraic $K$-group $G$, then $G$ is unirational if and only if $Z_{G}(T)$ is [Ros4, Prop.
 7.12].
@@
 In particular, we will see that unirationality of $U^{\operatorname{\rm{ab}}}$ implies unirationality of $U$ when $K$ has degree of imperfection $\leq 2$.
-We will also show that this result is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational but $U$ is not (though $U$ must be ext-unirational by Proposition 4.3).
+We will also show that this result is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational (though $U$ must be ext-unirational by Proposition 4.4).
 We first verify an analogous result for permawoundness in place of unirationality.
-The proof below makes reference to the notion of semiwoundness.
+For the definition of permawoundness, see either [Ros3, Def.
+1.2] or the beginning of §2.
+The proof below also makes reference to the notion of semiwoundness.
 Recall that a unipotent $K$-group scheme $U$ is called semiwound when it contains no copies of ${\mathbf{G}}_{\rm{a}}$ over $K$.
-This notion exhibits similar properties to the closely related notion of woundness, which is merely semiwoundness plus smoothness and connectedness; see [Ros2, Appendix A] for the basic properties of semiwound groups.
+This notion exhibits similar properties to the closely related notion of woundness, which is merely semiwoundness plus smoothness and connectedness; see [Ros3, Appendix A] for the basic properties of semiwound groups.
 Proposition 5.1.
@@
 Proof.
-All smooth unipotent groups are permawound over perfect fields [Ros2, Prop.
+All smooth unipotent groups are permawound over perfect fields [Ros3, Prop.
 5.2(i)], so we may assume that $K$ is imperfect.
@@
 We will show that $U$ is commutative, hence permawound by the already-treated commutative case.
-First we note that $U$ must be connected, as otherwise $U$ would admit a nontrivial étale unipotent quotient, hence a nontrivial commutative $p$-torsion étale quotient, which would imply that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is disconnected, in violation of [Ros2, Prop.
+First we note that $U$ must be connected, as otherwise $U$ would admit a nontrivial étale unipotent quotient, hence a nontrivial commutative $p$-torsion étale quotient, which would imply that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is disconnected, in violation of [Ros3, Prop.
 6.2].
@@
 The commutator map $U^{2}\rightarrow U$ descends to a map $\overline{U}^{2}\rightarrow U$, which further descends to a map $\overline{U}_{w}^{2}\rightarrow U$, where $\overline{U}_{w}$ is the maximal wound quotient of $\overline{U}$.
-This map is constant by [Ros3, Cor.
+This map is constant by [Ros4, Cor.
 10.3,Th.
@@
 Finally, let $U$ be an arbitrary smooth unipotent $K$-group scheme such that $U^{\operatorname{\rm{ab}}}/[p]U^{\operatorname{\rm{ab}}}$ is permawound, and let $U_{s}\trianglelefteq U$ be the maximal split $K$-subgroup.
-Then $U/U_{s}$ is permawound by the semiwound case, hence so is $U$, by Proposition 2.1 and [Ros2, Prop.
+Then $U/U_{s}$ is permawound by the semiwound case, hence so is $U$, by Proposition 2.1 and [Ros3, Prop.
 5.3].
@@
 Thanks to Proposition 5.1, we may assume that $U$ is commutative and $p$-torsion.
-When $K$ is perfect, the only wound group is the trivial group, and every smooth unipotent $K$-group is permawound [Ros2, Prop.
+When $K$ is perfect, the only wound group is the trivial group, and every smooth unipotent $K$-group is permawound [Ros3, Prop.
 5.2(i)], so we may assume that $K$ is imperfect.
-If $U_{s}\trianglelefteq U$ denotes the maximal split $K$-subgroup, then $U$ is permawound if and only if $U/U_{s}$ is, by Proposition 2.1 and [Ros2, Prop.
+If $U_{s}\trianglelefteq U$ denotes the maximal split $K$-subgroup, then $U$ is permawound if and only if $U/U_{s}$ is, by Proposition 2.1 and [Ros3, Prop.
 5.3].
 Since $U_{s}$ is killed under any $K_{s}$-homomorphism from $U$ into a wound group, we may assume $U$ is semiwound (that is, contains no copy of ${\mathbf{G}}_{\rm{a}}$).
-The only if direction now follows from [Ros3, Prop.
-10.1], plus the connectedness of $U$ (to ensure woundness of the maximal wound quotient) [Ros2, Prop.
+The only if direction now follows from [Ros4, Prop.
+10.1], plus the connectedness of $U$ (to ensure woundness of the maximal wound quotient) [Ros3, Prop.
 6.2].
 Now suppose that $K$ has finite nonzero degree of imperfection, and assume that $U$ is not permawound.
-We will show that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite, where (by abuse of notation) $\mathscr{V}$ is any $K$-form of the $K_{s}$-group which is denoted by $\mathscr{V}$ in [Ros2, Def.
+We will show that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite, where (by abuse of notation) $\mathscr{V}$ is any $K$-form of the $K_{s}$-group which is denoted by $\mathscr{V}$ in [Ros3, Def.
 7.3].
-By [Ros2, Th.
+By [Ros3, Th.
 1.4], we have for some $n>0$ (>0>0 because $U$ is not permawound) an exact sequence
@@
 We will show that $\operatorname{Ext}^{1}_{K_{s}}(W,\mathscr{V})$ is finite but $\operatorname{Ext}^{1}_{K_{s}}({\mathbf{G}}_{\rm{a}},\mathscr{V})$ is infinite, which will imply that $\operatorname{Hom}_{K_{s}}(U,\mathscr{V})$ is infinite as well.
-We may assume that $K=K_{s}$ [Ros3, Cor.
+We may assume that $K=K_{s}$ [Ros4, Cor.
 10.5].
@@
 We have a homomorphism $\phi\colon\operatorname{Ext}^{1}(W,\mathscr{V})\rightarrow\operatorname{Hom}(W,\mathscr{V})$ defined as follows: Given an extension $E$ in the former group, the Verschiebung endomorphism of $E$ descends to a map $W\rightarrow\mathscr{V}$ from the quotient $W$ to the subgroup $\mathscr{V}$.
-Then $\ker(\phi)$ consists of the extensions with trivial Verschiebung, and this is a trivial group by [Ros2, Cor.
+Then $\ker(\phi)$ consists of the extensions with trivial Verschiebung, and this is a trivial group by [Ros3, Cor.
 8.3].
@@
 We therefore assume that $U$ is semiwound.
-If $K$ has infinite degree of imperfection, then the assertion is immediate, thanks to [Ros2, Prop.
+If $K$ has infinite degree of imperfection, then the assertion is immediate, thanks to [Ros3, Prop.
 6.3].
 Thus we may assume that $K$ has finite degree of imperfection.
-Because permawound groups are unirational [Ros3, Th.
-1.9], we may use the rigidity property of permawound groups [Ros2, Th.
+Because permawound groups are unirational [Ros4, Th.
+1.9], we may use the rigidity property of permawound groups [Ros3, Th.
 1.5] and an easy induction to reduce to showing that any torsor for the $K$-group $V$ over a unirational $K$-scheme $X$ is also unirational, where $V$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$.
-But this assertion follows from [Ros3, Prop.
+But this assertion follows from [Ros4, Prop.
 8.6].
@@
 that sends a test scheme $T$ to the space of pointed $T$-morphisms $(X\times_{K}T,x_{T})\rightarrow(G\times_{K}T,1_{T})$.
-It turns out that it is represented by a smooth unipotent $K$-group scheme when $X$ is geometrically reduced of finite type and $G$ is wound unipotent, and in fact exhibits the somewhat stronger property that the corresponding scheme may be regarded as a subfunctor of the above morphism functor considered on the category of all $K$-schemes [Ros1, Th.
+It turns out that it is represented by a smooth unipotent $K$-group scheme when $X$ is geometrically reduced of finite type and $G$ is wound unipotent, and in fact exhibits the somewhat stronger property that the corresponding scheme may be regarded as a subfunctor of the above morphism functor considered on the category of all $K$-schemes [Ros2, Th.
 4.3].
@@
 We may assume that $K$ is separably closed.
-By [Ros3, Lem.
+By [Ros4, Lem.
 5.4], one may write $D_{i}$ as a finite disjoint union of $K$-schemes of the form $\mathrm{Spec}(K(\alpha_{ij}^{1/p^{n_{ij}}}))$ with $\alpha_{ij}\in K$, hence $\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{j}\operatorname{R}_{K(\alpha_{ij}^{1/p^{n_{ij}}})/K}({\mathbf{G}}_{\rm{m}})$, so we are reduced to the case in which $D_{i}=\mathrm{Spec}(K(\lambda_{i}^{1/p^{n_{i}}}))$ for all $i$.
 […]
-We proceed by induction on $r$, the case $r=0$ being trivial. Suppose that $r>0$. By the ubiquity property of permawound groups [Ros2, Th. 1.4], there is an exact sequence
+We proceed by induction on $r$, the case $r=0$ being trivial. Suppose that $r>0$. By the ubiquity property of permawound groups [Ros3, Th. 1.4], there is an exact sequence
 $0\longrightarrow U\longrightarrow W\longrightarrow{\mathbf{G}}_{\rm{a}}^{m}\longrightarrow 0$
@@
 In particular, $\dim(W)$ may be bounded depending only on $r$ and $d$.
-By the rigidity property of permawound groups [Ros2, Th.
+By the rigidity property of permawound groups [Ros3, Th.
 1.5], together with the fact that the number of terms in a filtration as in that result is bounded in terms of $\dim(W)$ and $p$, we may in fact assume that $U$ is either $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ or $\mathscr{V}$ (even though technically the former is not wound, being nonsmooth).
@@
 Thus we may assume that $\lambda_{1}\notin K^{p}$ and $n_{1}>0$.
-By [Ros3, Lem.
+By [Ros4, Lem.
 5.3], there is a totally nonsmooth (over $K$) $K$-subgroup $N\subset M:=\mathscr{M}\mathit{or}((\mathbf{P}^{1}\backslash\lambda_{1}^{1/p^{n_{1}}},\infty),(\mathscr{V},0))^{+}$ such that $(M_{K(\lambda_{1}^{1/p^{\infty}})})_{s}\subset N_{K(\lambda_{1}^{1/p^{\infty}})}$, where, for a unipotent $L$-group $V$, $V_{s}$ denotes the maximal split $L$-subgroup of $V$.
@@
 We first treat the special case in which the $X_{i}$ are all open subschemes of $\mathbf{P}^{1}_{K}$, in which case we prove a somewhat stronger result – namely, that the number of maps $f$ as in the proposition is bounded in terms only of $U$ and the degrees of the complementary divisors $\mathbf{P}^{1}\backslash X_{i}$.
-As in the proof of Lemma 5.4 above, we may invoke [Ros1, Th.
+As in the proof of Lemma 5.4 above, we may invoke [Ros2, Th.
 6.7] to conclude that it is sufficient to show that, for any reduced divisors $D_{1},\dots,D_{r}$ on $\mathbf{P}^{1}$, the number of multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ is finite and bounded in terms only of $U$ and the degrees of the $D_{i}$, and this is the content of Lemma 5.4(ii).
@@
 For technical reasons, we shall find it more convenient to prove a slightly modified version of the proposition.
-We invoke certain notation from the beginning of [Ros3, §7].
+We invoke certain notation from the beginning of [Ros4, §7].
 Namely, we have associated to $f$ a map $\Delta_{f,x_{1},\dots,x_{r}}\colon\prod_{i=1}^{r}X_{i}\rightarrow U,$ which enjoys two key properties: (1) $\Delta_{f,x_{1},\dots,x_{r}}=1$ whenever any coordinate is set to $x_{i}$, and (2) if $f$ vanishes whenever any coordinate is set to $x_{i}$, then $\Delta_{f,x_{1},\dots,x_{r}}=f$.
@@
 Theorem 5.7.
-Let $K$ be a field of degree of imperfection $r$, and let $U$ be a smooth connected unipotent $K$-group. Then $U$ is unirational if and only if $U/\mathscr{D}_{r}U$ is.
+Let $K$ be a field of degree of imperfection $r>1$. Then a smooth connected unipotent $K$-group $U$ is unirational if and only if $U/\mathscr{D}_{r}U$ is.
+For the corresponding result when $r=1$, see Corollary 4.5.
 Proof.
@@
 This descends to a map $c\colon\overline{U}^{r}\rightarrow U$, where $\overline{U}:=U/U^{\prime}$.
-Because $\mathscr{D}_{r}U\twoheadrightarrow\overline{U}/\mathscr{D}_{r}(\overline{U}), U¯\overline{U}$ is unirational by dimension induction.
+(Here we use $r>1$.) Because $\mathscr{D}_{r}U\twoheadrightarrow\overline{U}/\mathscr{D}_{r}(\overline{U}), U¯\overline{U}$ is unirational by dimension induction.
 If $c$ is constant, then $\mathscr{D}_{r}(U)=1$ and the theorem is trivially true.
@@
 Remark 5.8.
-If $U$ is a wound unirational $K$-group, and $K$ has degree of imperfection $r$, then $\mathscr{D}_{r+1}(U)=1$ [Ros3, Th.
+If $U$ is a wound unirational $K$-group, and $K$ has degree of imperfection $r$, then $\mathscr{D}_{r+1}(U)=1$ [Ros4, Th.
 1.5].
@@
 Of particular note in Theorem 5.7 is the case $r=2$, which says that if $K$ has degree of imperfection $2$, then $U$ is unirational if $U^{\operatorname{\rm{ab}}}$ is.
-We shall now give an example to show that this optimal, by constructing over every field of degree of imperfection $\geq 3$ a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not.
+The same holds for $r\leq 2$, by Corollary 4.5 and because every connected linear algebraic group is unirational when $r=0$.
+We shall now give an example to show that this is optimal, by constructing over every field of degree of imperfection $\geq 3$ a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational.
 Example 5.9.
@@
 That is, $b$ does not land in any proper $K$-subgroup scheme of $W$, or equivalently, for some $n>0$ the map $b^{n}\colon(V_{1,\lambda}\times V_{1,\mu})^{n}\rightarrow W$ obtained by adding $b$ over the $n$ coordinates is surjective.
-In fact, since all groups and maps live over the subfield $\mathbf{F}_{p}(\lambda,\mu)$, it suffices to verify this over this subfield, where it is [Ros3, Prop.
+In fact, since all groups and maps live over the subfield $\mathbf{F}_{p}(\lambda,\mu)$, it suffices to verify this over this subfield, where it is [Ros4, Prop.
 9.9].